### v1

# Large Angular Momentum

Abstract
Quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere. A spin $j$ (a $12j+$ 1-state system) for generic $j$ is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$. Here we study the angular momentum/spin in the limit, $j\to\infty$. The state, $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, where ${\hat{\mathbf{J}}}$ is the angular momentum operator and ${\mathbf{n}}$ stands for a generic unit vector in ${\bf R}^{3}$, is found to behave as a classical angular momentum, $j\,{\mathbf{n}}$. We discuss this phenomenon, by analysing the Stern-Gerlach experiments, the angular-momentum composition rule, and the rotation matrix. This problem arose from the consideration of a macroscopic body under an inhomogeneous magnetic field. Our observations help to explain how classical mechanics (with unique particle trajectories) emerges naturally from quantum mechanics in this context, and at the same time, make the widespread notion that large spins somehow become classical, a more precise one.
1 Introduction
It is a widely shared view that in the limit of large spin (angular momentum) the quantum-mechanical spin somehow becomes classical. After all, the large spin limit ($S/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ limit, and, in accordance with the general idea of the semi-classical limit, physics of large angular momentum is expected to become that of classical angular momentum. But to the best of the authors’ knowledge, it has not been clearly exhibited in the literature how exactly this works  [^1] Understandably, there is a vast literature about large angular momentum, from the early days of quantum mechanics to today, in a wide range of research fields, in various contexts and problems, and also based on different motivations. In order to avoid being biased and necessarily incomplete, we shall refrain in this work from attempting to cite or comment on any of them.. It is the aim of this note to fill in this gap.
The question came under a sharp focus in a general discussion on emergence of classical mechanics for a macroscopic body from quantum mechanics [1, 2]. Knowing about the typical Stern-Gerlach (SG) process for a spin $1/2$ atom, we ask what happens if a macroscopic body consisting of many (e.g. $N\sim\mathcal{O}(10^{23}))$ spins is set under an inhomogeneous magnetic field of a SG set-up. In general, with many spins oriented in different directions, the problem does not arise, as the bound-state Hamiltonian does not allow each component particle to split à la SG into spatially separated sub-wavepackets. Logically, however, one cannot exclude the special situation in which a large number of spins are all oriented in the same direction. Such a large spin system might continue to behave quantum-mechanically, e.g. with a typical SG spreading of the wavepackets, even in the $S/\hbar\to\infty$ limit.
It will be shown that the state $|S,{\mathbf{n}}\rangle$, the state in which the large spin is oriented in some direction ${\mathbf{n}}$, behaves instead as a classical angular momentum, yielding the unique trajectory in the SG processes (Sec. 2). This result leads us to consider more carefully in Sec. 3 the relation between the space of quantum-mechanical angular-momentum states and that of classical angular momenta, and to identify the special class of quantum angular momentum states to be put in correspondence with the classical angular momentum (see Eqs. (3.11), (3.12), (3.13) below). The content of the present work is to check the consistency of this picture, through the analysis of the SG experiments (Sec. 2), the angular-momentum composition rule (Sec. 4), and the rotation matrix (Sec. 5). We conclude with a brief reflection in Sec. 6.
2 Stern-Gerlach experiments
2.1 Spin $1/2$ atom
For definiteness let us take an incident atom with spin $S=\tfrac{1}{2}$ directed along a definite, but generic, direction ${\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, i.e.
$\Psi=\psi({\bf r},t)\,|{\bf n}\rangle\;,$ (2.1)
$|{\bf n}\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,\qquad c_{1}=e^{-i\phi/2}\cos\tfrac{\theta}{2}\;,\qquad c_{2}=e^{i\phi/2}\sin\tfrac{\theta}{2}\;.$ (2.2)
When the atom passes through a region with an inhomogeneous magnetic field,
${\mathbf{B}}=(0,0,B(z)),\qquad\frac{\partial B(z)}{\partial z}\neq 0\;,$ (2.3)
the wave packet splits into two subpackets, with respective weights, $|c_{1}|^{2}$ and $|c_{2}|^{2}$: they manifest as the relative intensities of the image bands on the screen at the end.
The original Stern-Gerlach experiment [3] has had a fundamental impact to the development of quantum theory, in proving the quantization of the angular momentum, the existence of half-integer spin, and more generally, demonstrating the quantum-mechanical nature of the center of mass (CM) of the silver atom [2].
2.2 Large spin
A classical particle with a magnetic moment directed towards
${\boldsymbol{\mu}}={\rm const}.\,{\bf n}\;,\qquad{\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;,$ (2.4)
moves in an inhomogeneous magnetic field (2.3), according to Newton’s equations,
$m\,{\dot{\bf r}}={\bf p}\;,\qquad\frac{dp_{x}}{dt}=\frac{dp_{y}}{dt}=0\;,\quad\frac{dp_{z}}{dt}=F_{z}=\frac{\partial}{\partial z}{\boldsymbol{\mu}}\cdot{\bf B}\;.$ (2.5)
The way the unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [1], where the magnetic moment is the expectation value
$\sum_{i}\langle\Psi|({\hat{\boldsymbol{\mu}}}_{i}+\frac{e_{i}{\hat{\boldsymbol{\ell}}}_{i}}{2m_{i}c})|\Psi\rangle={\boldsymbol{\mu}}\;,$ (2.6)
taken in the internal bound-state wave function $|\Psi\rangle$ where $\mu_{i}$ and $\frac{e_{i}{\bf\ell}_{i}}{2m_{i}c}$ denote the intrinsic magnetic moment and the one due to the orbital motion of the $i$-th constituent atom or molecule ($i=1,2,\ldots,N$), respectively. Clearly, in general, the considerations made in Sec. 2.1 for a spin $1/2$ atom, with a doubly split wavepacket, cannot be simply generalized to (or compared with) a classical body (2.6) with $N\sim\mathcal{O}(10^{23})$ constituents.
Typically, the spins are oriented along different directions, but the particles carrying them are bound in the macroscopic body made of atomic, molecular and larger structures. The wave function of particles in bound states does not split in sub-wavepackets, as that of an isolated spin $\tfrac{1}{2}$ atom does, under an inhomogeneous magnetic field. The bound-state Hamiltonian does not allow it  [^2] This discussion is indeed closely parallel to the question of quantum diffusion. Unlike free particles, particles in bound states (the electrons in atoms; atoms in molecules, etc.) do not diffuse, as they move in binding potentials. This is one of the elements for the emergence of the classical mechanics, with unique trajectories for macroscopic bodies. As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is suppressed by mass.. What determines the motion of the CM of the body is the average magnetic moment, (2.6). Logically, nevertheless, one cannot exclude special cases (e.g. a magnetized piece of metal) of all spins of the body oriented in the same direction. One might wonder how the wave function of such a particle with spin $S\to\infty$ behaves, under an inhomogeneous magnetic field.
The question is whether the conditions enlisted in Ref. [1] (see Appendix A), for the emergence of classical mechanics for a macroscopic body with a unique trajectory, are sufficient, or whether some extra condition (or a new mechanism) is required, to suppress possible wide spreading of the wave function into many sub-packets (see Fig. 1 for a small spin, $S=13/2$) under an inhomogeneous magnetic field.
Figure 1: The distribution $|c_{k}|^{2}$ in $k$, i.e. in possible values of $S_{z}$, $-S\leq S_{z}\leq S$, for a spin $S=13/2$ particle in the state (2.7), with $\theta=\pi/2$ (center, orange dots) and $\theta=\pi/4$ (right, blue dots).
The answer turns out to be simple but somewhat unexpected. The following discussion, up to Eq. (2.15), is taken almost verbatim from Sec. 3.1.3 of Ref. [2]; for the present work it represents the starting point, and serves as the main tool of analysis below.
Consider the state of a spin $S$ directed towards a direction ${\mathbf{n}}$,
$\left({\hat{\mathbf{S}}}\cdot{\mathbf{n}}\right)\,|S,{\mathbf{n}}\rangle=S\,|S,{\mathbf{n}}\rangle\;,\qquad{\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (2.7)
Choosing the magnetic field (and its gradients) in the ${z}$-direction, we need to express $|S,{\mathbf{n}}\rangle$ as a superposition of the eigenstates of ${\hat{S_{z}}}$ ($S=N/2$),
$|S,{\mathbf{n}}\rangle=\sum_{k=0}^{N}c_{k}\,|S,S_{z}=M\rangle\;,\qquad M=-\frac{N}{2}+k\;,\quad(k=0,1,\ldots,N)\;.$ (2.8)
The expansion coefficients $c_{k}$ are known:
$c_{k}={\binom{N}{k}}^{1/2}\,e^{-iM\phi/2}\left(\cos\tfrac{\theta}{2}\right)^{k}\left(\sin\tfrac{\theta}{2}\right)^{N-k}\;,\qquad\sum_{k=0}^{N}|c_{k}|^{2}=1\;,$ (2.9)
where $\binom{N}{k}=N!/k!(N-k)!$ are the binomial coefficients.
To get (2.9), let us consider (2.7) as a direct product of $N$ spin $\tfrac{1}{2}$ particles, (2.2),
$|S,{\mathbf{n}}\rangle=|{\bf n}\rangle_{1}\otimes|{\bf n}\rangle_{2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (2.10)
all oriented in the same direction. Expanding and collecting terms with a fixed $k$ (the number of spin-up particles) one arrives at Eq. (2.9).
Using Stirling’s formula, one can derive, for large values of both $N$ and $k$ with $x=k/N$ fixed, the distribution across different values of $S_{z}=M$,
$|c_{k}|^{2}\simeq e^{Nf(x)}\;,\qquad x=k/N\;,$ (2.11)
with
$f(x)=-x\log x-(1-x)\log(1-x)+2x\log\cos\tfrac{\theta}{2}+2(1-x)\log\sin\tfrac{\theta}{2}\;.$ (2.12)
The saddle-point approximation, valid at $N\to\infty$, yields
$f(x)\simeq-\frac{(x-x_{0})^{2}}{x_{0}(1-x_{0})}\;,\qquad x_{0}=\cos^{2}\tfrac{\theta}{2}\;,$ (2.13)
and thus
[…]
in the $N\to\infty$ ($x=k/N$ fixed) limit. The narrow peak position $x=x_{0}$ corresponds to (see Eq. (2.8))
$S_{z}=M=N(x-\tfrac{1}{2})=S\,(2\cos^{2}\tfrac{\theta}{2}-1)=S\,\cos\theta\;.$ (2.15)
This means that a large spin ($S\gg\hbar$) quantum particle with spin directed towards ${\mathbf{n}}$, in a Stern-Gerlach setting with an inhomogeneous magnetic field (2.3), moves along a single trajectory of a classical particle with $S_{z}=S\,\cos\theta$, instead of spreading over a wide range of split sub-packet trajectories covering $-S\leq S_{z}\leq S$. See Fig. 2, cfr. Fig. 1.
Figure 2: The distribution $|c_{k}|^{2}$ in $k$, i.e. in possible values of $S_{z}$, for a spin $S=1\cdot 10^{3}$ particle in the state (2.7), with $\theta=\pi/2$ (center, orange peak) and $\theta=\pi/4$ (right, blue peak). The particle starts looking like classical.
This (perhaps) somewhat surprising result shows that quantum mechanics (QM) takes care of itself, so to speak, in ensuring that a large spin particle ($S/\hbar\to\infty$) traces a classical trajectory. This is consistent with the known general behavior of the wave function in the semi-classical limit ($\hbar\to 0$)  [^3] Of course this does not mean that the classical limit necessarily requires or implies $S/\hbar\to\infty$, though.. Note that if the value $S\gg\hbar$ is understood as due to the large number of spin $1/2$ particles composing it (see the comment after (2.9)), the spike (2.13), (2.14) can be understood as due to the accumulation of a huge number of microstates giving $S_{z}\simeq M$.
Another, equivalent way to see the shrinkage of the distribution in $S_{z}$, (2.14), is to study its dispersion. Given the relative weights $|c_{k}|^{2}$, (2.9), where $S_{z}=-S+k$ and $S=\tfrac{N}{2}$, with $k=0,1,\ldots,N$, it is a simple exercise to find that
$\langle S_{z}\rangle=S\,\cos\theta\;,\qquad\langle S_{z}^{2}\rangle=S^{2}\cos^{2}\theta+\frac{S}{2}\sin^{2}\theta\;,$ (2.16)
so that the dispersion (the fluctuation width) is given by
$(\Delta S_{z})^{2}=\langle(S_{z}-\langle S_{z}\rangle)^{2}\rangle=\langle S_{z}^{2}\rangle-\langle S_{z}\rangle^{2}=\frac{S}{2}\sin^{2}\theta\;.$ (2.17)
It follows that
$\lim_{S/\hbar\to\infty}\frac{|\Delta S_{z}|}{S}=0\;,$ (2.18)
i.e. it is an infinitely narrow Gaussian. Note that the SG set-up is indeed a measurement of $S_{z}$ in the state (2.7). The fact that the result is always $S_{z}\sim\langle S_{z}\rangle=S\,\cos\theta$ with no dispersion, means that it is a classical angular momentum $\sim S{\mathbf{n}}$ (see Eqs. (3.11), (3.12) below).
2.2.1 Density matrix
The same phenomenon can be seen also by using the density matrix. Being the state (2.7) a pure state, its $(2S+1)\times(2S+1)$ density matrix is given by
$\rho_{k\ell}=c_{k}c_{\ell}^{*}\,,$ (2.19)
where $c_{k}(c_{\ell})$ are given in Eq. (2.9). In the limit as $S\to\infty$, it can be observed that it essentially reduces to a single diagonal element, provided that the fluctuations of the order of $\sqrt{S}$ are neglected in comparison to $S$. The density matrix ${\boldsymbol{\rho}}$ for $S=500$, with $(\theta,\phi)=(\frac{\pi}{2},0)$ and $(\theta,\phi)=(\frac{\pi}{4},0)$, is depicted in Figs. 3.
Figure 3: The density matrix ${\boldsymbol{\rho}}$ for the state (2.7), of spin $S=5\cdot 10^{2}$, with ($a$): $(\theta,\phi)=(\tfrac{\pi}{2},0)$ and ($b$): $(\theta,\phi)=(\tfrac{\pi}{4},0)$. The height of the plots indicates the matrix elements $\rho_{k\ell}$. Note that ${\boldsymbol{\rho}}$ is real for $\phi=0$.
2.2.2 The state preparation and successive SG set-ups
For a small $S$, as it is well known, the SG set-up can be used to prepare any desired state $|S,S_{z}=M\rangle$, by an intentionally biased SG apparatus, e.g. in which all the split sub-wavepackets are blocked, except for the one corresponding to $|S,S_{z}=M\rangle$. Equivalently, one may use the so-called null-measurements (see Ref. [4] for a recent review). In any case, for a small finite $S$, where there are $12S+$ 1 well-separated sub-packets (see Fig. 1), there is no difficulty in extracting any desired quantum state for subsequent experiments (state preparation).
In the large-spin limit ($S\to\infty$) we are considering, since there is essentially only one wavepacket present (see (2.15)), there is no possibility of extracting or selecting the desired sub-wavepacket. The experimental resolution would not allow it. We have in mind a macroscopic scenario where, in comparison to the finite region (in principle) occupied by the waves $-S\leq M\leq S$, the fluctuations of the values of $M$ around (2.15) are of the order of $\sqrt{S}\ll S$, i.e. of zero width.
We might consider introducing a second set of Stern-Gerlach (SG) magnetic fields directed in a different direction, denoted by ${\hat{z}}^{\prime}$, as a new quantization axis. This setup allows us to study the state emerging from the first set of SG magnetic fields (without the presence of an imaging screen),
$|S,{\mathbf{n}}\rangle\sim c_{k}\,|S,S_{z}=M\rangle\;,\qquad k\sim 2S\cos^{2}\tfrac{\theta}{2}\;,\qquad M\sim S\,\cos\theta\;.$ (2.20)
Since the sum of states around $k\sim 2S\cos^{2}\left(\frac{\theta}{2}\right)$ is essentially equal to the complete sum, this state can be rewritten in terms of the eigenvalues $M^{\prime}=-S+k^{\prime}$, where $k^{\prime}=0,1,\ldots 2S$ correspond to the eigenvalues of $S_{z^{\prime}}$ as:
$|S,{\mathbf{n}}\rangle$ $=$ $\sum_{k,k^{\prime}}c_{k}\,|S,M^{\prime}\rangle\langle S,M^{\prime}|S,M\rangle$ (2.21)
$=$ $\sum_{k,k^{\prime}}|S,M^{\prime}\rangle\langle S,M^{\prime}|S,M\rangle\langle S,M|S,{\mathbf{n}}\rangle$
$=$ $\sum_{k^{\prime}}|S,M^{\prime}\rangle\langle S,M^{\prime}|S,{\mathbf{n}}\rangle=\sum_{k^{\prime}}c_{k^{\prime}}|S,M^{\prime}\rangle\;.$
This represents an expansion of the original state $|S,{\mathbf{n}}\rangle$ as a linear combination of the eigenstates of $S_{z^{\prime}}$. In the limit $S\to\infty$, the results from Section 2.2 tell us that the sum is dominated by the small region around $S_{z^{\prime}}\sim S,\cos({\mathbf{n}}\cdot{\hat{z}}^{\prime})$, with negligible fluctuations. The trajectory remains unique, regardless of the choice of quantization axis (i.e. the direction of the magnetic field) ${\hat{z}}^{\prime}$.
2.2.3 States almost directed towards ${\vec{\mathbf{n}}}$
Obviously, in the large spin limit, other states close to $|S,{\mathbf{n}}\rangle$,
$\left({\hat{\mathbf{S}}}\cdot{\mathbf{n}}\right)\,|S,{\mathbf{n}},q\rangle=(S-q)\,|S,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,\,\,{\rm fixed}$ (2.22)
(i.e. the spin roughly oriented towards ${\mathbf{n}}$) will behave similarly. For instance, the $q=1$ state is (cfr. Eq. (2.10))
$|S,{\mathbf{n}},1\rangle$ $=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}\,|{\bf n}\rangle_{1}\otimes\cdots|{\bf n}\rangle_{\ell}\otimes|-{\bf n}\rangle\otimes|{\bf n}\rangle_{\ell+2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (2.23)
$=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}|S_{1},{\mathbf{n}}\rangle\otimes|-{\bf n}\rangle\otimes|S_{2},{\mathbf{n}}\rangle\;,$
where $|S_{1},{\mathbf{n}}\rangle$ and $|S_{2},{\mathbf{n}}\rangle$ ($S_{1}=\tfrac{\ell}{2}\;,S_{2}=\tfrac{N-\ell-1}{2}$), are two spins oriented towards ${\mathbf{n}}$, (2.8), as already studied in Sec. 2.2. In the $S\to\infty$ limit, the sum is dominated by terms with $\ell\sim\mathcal{O}(N)\to\infty$. The SG projection of this state onto the $S_{z}$ eigenstates is therefore approximated by
$S_{z}\sim S_{1\,z}+S_{2\,z}=S_{1}\cos\theta+S_{2}\cos\theta\simeq S\,\cos\theta\;,$ (2.24)
where the results of Sec. 2.2 have been used. A similar result holds for $|S,{\mathbf{n}},q\rangle$ for any finite $q$, showing that all these states behave as a classical angular momentum, $S\,{\mathbf{n}}$.
3 Quantum and classical angular momentum states
Before proceeding, we must discuss the correspondence between the quantum mechanical angular momentum states and their classical counterparts, more carefully. The spin $1/2$ (two-state system) wave function is given by
$|\psi\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,$ (3.1)
where $(c_{1},c_{2})\sim\lambda(c_{1},c_{2})$ with $\lambda\in{\mathbf{C}}/\{0\}$, describes the points of $\mathbf{CP}^{1}\sim S^{2}$. We note that for a spin $1/2$, any pure state (3.1) can be interpreted as the state in which the spin is oriented in some space direction ${\mathbf{n}}$, as in Eq. (2.2), without loss of generality. This is because the manifold of pure spin $1/2$ quantum states and that of the unit space vector are both $S^{2}\sim\mathbf{CP}^{1}$. Indeed, given the state (2.2), there is always a rotation matrix such that
$D^{1/2}(0,\theta,\phi)\cdot\left(\begin{array}[]{c}e^{-i\phi/2}\cos\tfrac{\theta}{2}\\ e^{i\phi/2}\sin\tfrac{\theta}{2}\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\;.$ (3.2)
We recall that the general rotation matrix for a spin $1/2$ is given, using Euler angles  [^4] Note that the third Euler angle (the rotation angle $\gamma$ about the final $z$-axis) is redundant here, as it gives only the phase $e^{i\gamma/2}$, and has been set to 00 in (3.2). ($\gamma,\theta,\phi$), as follows
$D^{1/2}(\gamma,\theta,\phi)=U_{z}(\gamma)U_{y}(\theta)U_{z}(\phi)=\left(\begin{array}[]{cc}e^{i(\gamma+\phi)/2}\cos\tfrac{\theta}{2}&e^{(\gamma-\phi)/2}\sin\tfrac{\theta}{2}\\ -e^{i(-\gamma+\phi)/2}\sin\tfrac{\theta}{2}&e^{-i(\gamma+\phi)/2}\cos\tfrac{\theta}{2}\end{array}\right)\;.$ (3.3)
The situation is different for a generic spin $j=N/2$  [^5] In this section we denote the general angular momentum (including orbital angular momentum) by $j$, rather than $S$, a symbol adopted commonly to indicate the spin (the intrinsic angular momentum of a particle). In this work we will use both $j$ and $S$. Likewise, we shall use the terms “spin” and “angular momentum”, interchangeably.. The wave function now has the form,
$|\psi\rangle=\sum_{k=0}^{N}b_{k}|j,j_{z}=-j+k\rangle$ (3.4)
where
$(b_{0},b_{1},\ldots,b_{N})\sim\lambda(b_{0},b_{1},\ldots,b_{N})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (3.5)
are the coordinates of the space $\mathbf{CP}^{N}$. On the other hand, the variety of the directions in the three-dimensional space, ${\mathbf{n}}$, is always $S^{2}\sim\mathbf{CP}^{1}$. It means that the state of the spin oriented towards a definite direction (2.8), is not a general pure spin $j$ state, but belongs to a small subset, $\mathbf{CP}^{1}\subset\mathbf{CP}^{2j}$. In other words, not all states of spin $j$, as described in (3.4), can be transformed by a rotation matrix (selecting an appropriate new ${\hat{z}}$ axis) into the form:
$(1,0,\ldots,0)^{T}\;.$ (3.6)
These considerations highlight a potential subtlety in our argument that quantum mechanical angular momentum states somehow become classical in the large $j$ limit.
Let us recall that because of the commutation relations
$[{\hat{J}}_{k},{\hat{J}}_{\ell}]=i\hbar\,\epsilon^{k\ell m}{\hat{J}}_{m}\;,$ (3.7)
only one of the components, e.g. ${\hat{J}}_{z}$, can be diagonalized together with the total (Casimir) angular momentum, ${\hat{J}}^{2}={\hat{J}}_{x}^{2}+{\hat{J}}_{y}^{2}+{\hat{J}}_{z}^{2}$,
${\hat{J}}^{2}|j,m\rangle=j(j+1)|j,m\rangle\;,\qquad{\hat{J}}_{z}|j,m\rangle=m|j,m\rangle\;.$ (3.8)
In the state, $|j,m\rangle$, where ${\hat{J}}_{z}$ has a definite value, $m\,\hbar$, ${\hat{J}}_{x}$ and ${\hat{J}}_{y}$ are fluctuating. Their magnitudes $\Delta{J}_{x,y}=\langle j,m|{\hat{J}}_{x,y}^{2}|j,m\rangle^{1/2}$ are found to be
$\Delta{J}_{x}=\Delta{J}_{y}=\left(\langle j,m|\big{(}{\hat{J}}^{2}-{\hat{J}}_{z}^{2}\big{)}|j,m\rangle/2\right)^{1/2}=\left(j(j+1)-m^{2})/2\right)^{1/2}\hbar\sim\mathcal{O}(j)\hbar\;.$ (3.9)
Therefore, in a generic state $|j,m\rangle$, there are strong quantum fluctuations of $J_{x}$ and $J_{y}$ at $j\to\infty$. There is no way such a state can be associated with a classical angular momentum, which requires all three components to be well-defined simultaneously.
The exception occurs for states $|j,\pm j\rangle$, for which
$\Delta{\hat{J}}_{x}=\Delta{\hat{J}}_{y}=(j/2)^{1/2}\hbar\ll j\,\hbar\;.$ (3.10)
For these state, it makes sense to interpret them “classically” as angular momentum pointed towards ±$\pm{\hat{z}}$, as its fluctuations $(\sim\mathcal{O}(\sqrt{j})$) become negligible with respect to its magnitude $j$, in the limit $j\to\infty$. The same holds true in all states in which spin is oriented towards a generic direction ${\mathbf{n}}$, as has been already noted in the previous section, Eqs. (2.16)-(2.18). Accordingly, we propose the correspondence (by setting $\hbar=1$)
$\boxed{\quad|j,{\mathbf{n}}\rangle\quad\longleftrightarrow\quad{\bf j}^{(\rm class)}=j\,{\bf n}\;\quad}$ (3.11)
where
${\hat{J}}^{2}|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\;.$ (3.12)
Actually, all states “almost oriented towards ${\mathbf{n}}$”, $|j,{\mathbf{n}},q\rangle,$
$\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,,\,\,\,{\rm finite},$ (3.13)
discussed in Sec. 2.2.3 (see (2.22)), can also be interpreted as ${\bf j}^{(\rm class)}=j\,{\bf n}$. The aim of our present investigation is to examine the consistency of such a picture, with respect to the SG processes (this has already been done in Sec. 2.2), the addition rule of the angular momenta (Sec. 4 below) and the rotation matrix (Sec. 5 below).
As for the states with strong quantum fluctuations $|j,m\rangle$, (3.9) (with $m$ generic and not close to ±$\pm j$), they are certainly quantum mechanical, independently of $j$. The same can be said of a generic spin $j$ wave function, (3.4). If such a particle is sent into a SG set-up, it could split in an arbitrary distribution of sub wavepackets, in contrast to the cases discussed in Sec. 2.2. This results in a macroscopic quantum system. It is possible that such a strongly fluctuating, macroscopic quantum angular-momentum state can be realized experimentally, if the whole system is brought to extremely low temperatures near $T=0$  [^6] We recall that any system is in its quantum-mechanical ground state at $T=0$. Accordingly many macroscopic systems can be brought to a quantum-mechanical state, if sufficiently low temperatures can be attained. See the references cited in [1] for the recent efforts to fabricate such systems experimentally..
To see what happens at finite temperatures (either the body-temperature of the macroscopic body, or a warm environment) instead, it is useful to recall the so-called Born-Einstein dispute and its resolution (see [1]). Einstein strongly rebuked the idea by Born, that the absence of quantum diffusion is sufficient to explain the classical nature (the unique trajectoriy) of a macroscopic body, arguing that a doubly or multiply split wave packet, with their centers separated by macroscopic distances, are allowed by Schrödinger equations even for a macroscopic particle. The latter is certainly non-classical. The missing piece for solving this puzzle turns out to be the temperature [1]. Even though at exceptionally low temperatures such a state is possible, it is not so at finite temperatures. Emission of photons and the ensuing self-decoherence (in the case of an isolated body) [1] or an environment-induced decoherence (for an open system) [5, 6, 7] makes a split system a mixture (a mixed state) essentially instantaneously. Also, they cannot be prepared experimentally, e.g., by a passage through a double slit [1]. A (macroscopic) particle passes either through one or the other slit, due to absence of diffusion.
A similar consideration applies to the general angular momentum states with strong fluctuations, (3.9) or (3.4). In the large $j$ limit, such a state necessarily involves large space extension: a macroscopic quantum state. A macroscopic pure quantum state can be realized only at exceedingly low temperatures  [^7] To tie up possible loose ends of the argument, let us note that a single, spin $1/2$ isolated atom in its ground state (a microscopic system) travelling in a good vacuum, has effectively $T=0$. Thus the fact that its split wave packets can attain a macroscopic extension - i.e. its being in a macroscopic quantum state - is perfectly consistent. To realize a macroscopic system (made of many atoms and molecules) in a pure quantum state is another story: it is much more difficult to prepare the necessary low temperatures close to $T=0$. . See [1] for more discussions. At finite temperatures, macroscopically split wave packet of such a particle under the SG set-up is necessarily an incoherent mixture.
What is rather remarkable, perhaps, is the fact that the angular momentum states of minimum fluctuations, (3.12), (3.13), though pure, effectively become classical at $j\to\infty$, without need of any (thermal or environment-induced) decoherence effects, hence independently of the temperatures. This is in agreement with the known general semi-classical behavior of a (pure-state) wave function in QM in the limit $\hbar\to 0$.
3.1 Orbital angular momentum
All our formal arguments (3.7)–(3.13) apply equally well, both to spin (the intrinsic angular momentum of a particle or the total angular momentum of the center of mass of a body in its rest frame) and to orbital angular momentum. However, our primary tool of analysis is the Stern-Gerlach (SG) set-up (Sec. 2.2), which is more suitable for spin than for an angular momentum associated with the orbital motion of a particle. One might wonder how the large angular momentum limit works for the latter. Here is a brief acccount.
The (angular part of the) wave function of a particle in the state $|\ell,m\rangle$ corresponds to the spherical harmonics,
$\langle\theta,\phi|\ell,m\rangle=Y_{\ell,m}(\theta,\phi)=(-)^{(m+|m|)}\sqrt{\tfrac{(2\ell+1)(\ell-|m|)!}{4\pi(\ell+|m|)!}}P_{\ell}^{|m|}(x)\,e^{im\phi}\;,\qquad x\equiv\cos\theta\;,$ (3.14)
where $P_{\ell}^{m}(x)$ are the associated Legendre polynomials. Let us consider the state of minimum fluctuation, $|\ell,\ell\rangle$ (see Eq. (3.10)). It is easy to see that the angular distribution is strongly peaked at $\theta=\tfrac{\pi}{2}$ at large $\ell$,
$|Y_{\ell,\ell}(\theta,\phi)|^{2}=\frac{2\ell+1}{4\pi}\frac{(2\ell)!}{2^{2\ell}(l!)^{2}}(1-x^{2})^{\ell}\stackrel{{\scriptstyle\ell\to\infty}}{{\longrightarrow}}\frac{1}{2\pi}\delta(x)\;.$ (3.15)
For instance, see Fig. 5 for $\ell=m=10^{3}$. In states $|\ell,m\rangle$ with stronger fluctuations ($m$ not close to $\ell$), the distribution instead covers a large portion of the $3$D solid angle, for any $\ell$ (see Fig. 5 for $\ell=100,m=50$).
Now a classical angular momentum has its all three components well defined. By choosing an appropriate coordinate system, it can be written as ${\bf L}^{(\rm class)}=(0,0,L)$. It describes a particle moving in the $(x,y)$ plane, with $z=p_{z}=0$. The large $\ell$ limit of the wave function $|Y_{\ell,\ell}(\theta,\phi)|^{2}$, (3.15), describes precisely such a motion.
Figure 4: The distribution in $\theta$, for
the state $|\ell,\ell\rangle$ with $\ell=10^{3}$.
Figure 5: The distribution in $\theta$, for
the state $|\ell,m\rangle$ with $\ell=100$, $m=50$.
4 Addition of angular momenta
Let us now consider two spins, ${\mathbf{j}}_{1}$ and ${\mathbf{j}}_{2}$. The composition-decomposition rules in QM is well known, i.e.
${\underline{j_{1}}}\otimes{\underline{j_{2}}}={\underline{j_{1}+j_{2}}}\oplus{\underline{j_{1}+j_{2}-1}}\oplus\ldots{\underline{{|j_{1}-j_{2}|}}}\;,$ (4.1)
where $\underline{j_{1}}$, $\underline{j_{2}}$, indicates the multiplet (the irreducible representation) of SU(2)  [^8] We recall that SU(2) is the simply-connected double cover of the orthogonal group SO(3)., with multiplicity $12j+$ 1 and with $\max\{j_{z}\}=j$. We set $\hbar=1$ here.
We wish to find out how the addition rule looks like in the limit $j_{1},j_{2}\to\infty$. For the reasons discussed in Sec. 3, we are particularly interested in the composition rule for angular momentum states with minimum fluctuations, of the form, $|j,{\mathbf{n}}\rangle$. Therefore, let us consider two particles (spins) in the states,
${\hat{\mathbf{j}}}_{1}\cdot{\mathbf{n}}_{1}\,|j_{1},{\mathbf{n}}_{1}\rangle=j_{1}\,|j_{1},{\mathbf{n}}_{1}\rangle\;,\qquad{\hat{\mathbf{j}}}_{2}\cdot{\mathbf{n}}_{2}\,|j_{2},{\mathbf{n}}_{2}\rangle=j_{2}\,|j_{2},{\mathbf{n}}_{2}\rangle\;,$ (4.2)
namely, they are spins oriented towards the directions, ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$, respectively. Our aim is to find out the properties of the direct product state,
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\;,$ (4.3)
in the large $j_{1},j_{2}$ limit. Since the choice of the axes is arbitrary, we may take
${\mathbf{n}}_{2}=(0,0,1)\;.$ (4.4)
Then the product state is just
$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle\otimes|j_{2},m_{2}=j_{2}\rangle\;,$ (4.5)
where $|j_{2},m_{2}=j_{2}\rangle$ is the highest $\hat{j}_{2\,z}$ state. We added the tilde sign on ${\mathbf{n}}_{1}$ to indicate that it is the vector ${\mathbf{n}}_{1}$ in the reference system of (4.4). Now $|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle$ in the $j_{1}\to\infty$ limit has been studied in Sec. 2. It is
$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle=\sum_{m=-j_{1}}^{j_{1}}c_{m}\,|j_{1\,z}=m\rangle\;,\qquad$ (4.6)
where the sum is dominated by the values of $m$ around
$m=j_{1\,z}\simeq j_{1}\cos\theta\;,\qquad\cos\theta={\tilde{\mathbf{n}}}_{1\,z}={\mathbf{n}}_{1}\cdot{\mathbf{n}}_{2}\;.$ (4.7)
Since the eigenvalues of $\hat{j_{z}}$ simply add up in the product state, the expansion of $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle$ in the eigenstates of $\hat{j}_{{\rm tot}\,z}$ is dominated by terms
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{{\tilde{n}}_{2}}\simeq j_{1}\cos\theta+j_{2}\rangle\;.$ (4.8)
Exchanging ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$ and repeating the arguments, one finds also that
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{\tilde{n}_{1}}\simeq j_{1}+j_{2}\cos\theta\rangle\;.$ (4.9)
But which quantum angular-momentum state does the product state (4.3) represent? The answer is that
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},{\mathbf{n}}\rangle\;,$ (4.10)
where $j_{\rm tot}$ and the unit vector ${\mathbf{n}}$ are such that (i.e. defined by)
$j_{\rm tot}\,{\mathbf{n}}=j_{1}{\mathbf{n}}_{1}+j_{2}{\mathbf{n}}_{2}\;.$ (4.11)
The proof of this is as follows. By selecting the directions of the SG magnets directions (the quantization axis) to ${\mathbf{n}}_{2}$ or to ${\mathbf{n}}_{1}$ on the state $|j_{\rm tot},{\mathbf{n}}\rangle$, (4.11), one can apply the results of Sec. 2.1 to get exactly the results (4.8) or (4.9), respectively. Now, according to the quantum-classical correspondence (Eq. (3.11) and Eq. (3.12)), the state $|j_{\rm tot},{\mathbf{n}}\rangle$ (see Eq. (4.11)) translates to
${\mathbf{j}}_{\rm tot}^{(\rm class)}=j_{1}\,{\mathbf{n}}_{1}+j_{2}\,{\mathbf{n}}_{2}\;.$ (4.12)
But, in view of (3.11), this is precisely the addition rule of the two classical angular momenta (i.e. a vector sum).
To be complete, one must check that the projection on any generic direction ${\mathbf{n}}_{3}$ works correctly. By using the result of Sec. 2 for the SG projection on a new ${\mathbf{n}}_{3}$ axis, one finds
$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot}^{\prime},j_{\rm tot}^{\prime\,n_{3}}\simeq j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}\rangle\;,$ (4.13)
where we made use of the fact that ${\hat{\mathbf{j}}}_{1}$ and ${\hat{\mathbf{j}}}_{2}$ commute and the eigenvalues of their ${\mathbf{n}}_{3}$ components just sum up in the direct product state. Also, to be conservative, we have left $j_{\rm tot}^{\prime}$ unknown. However,
$j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}=(j_{1}{\mathbf{n}}_{1}+j_{2}{\mathbf{n}}_{2})\cdot{\mathbf{n}}_{3}\;,$ (4.14)
thus
${\mathbf{j}}_{\rm tot}^{\prime}={\mathbf{j}}_{\rm tot}=j_{1}\,{\mathbf{n}}_{1}+j_{2}\,{\mathbf{n}}_{2}\;,$ (4.15)
by again using the results of Sec. 2.1 in Eq. (4.13). In conclusion, the direct product state, (4.2), (4.3), has a simple classical interpretation. It corresponds to the classical sum (the vector addition) of two angular momenta.
From the perspective of the general composition rule of two angular momentum states (4.1), what we have seen are the results concerning a specific pair of states (i.e. those of minimum fluctuations) in ${\underline{j_{1}}}$ and ${\underline{j_{2}}}$, characterized by the “orientations” ${\bf n}_{1}$ and ${\bf n}_{2}$. The total angular momentum magnitude in the product state (4.11) satisfies
$j_{1}+j_{2}\geq j_{\rm tot}=|j_{1}{\mathbf{n}}_{1}+j_{2}{\mathbf{n}}_{2}|\geq|j_{1}-j_{2}|\;,$ (4.16)
depending on the relative orientation of ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$. This is consistent with the quantum-mechanical addition law, (4.1).
5 Rotation matrix
The rotation matrix for a general spin $j$ can be constructed by taking the direct products of the rotation matrix for spin $1/2$, i.e. (3.3). This is based on the fact that any angular momentum state, $|j,m\rangle$, can be constructed as the system made of $N$ spin $\tfrac{1}{2}$ particles, completely symmetric under exchange of the particles, a fact already conveniently used in Sec. 2.1. The theory of angular momentum can be fully reconstructed this way [8] by using the operators,
${\hat{J}}^{a}=\sum_{i,j=1,2}a_{i}^{\dagger}\Big{(}\frac{\tau^{a}}{2}\Big{)}_{ij}a_{j}\;,\qquad[{\hat{J}}^{a},{\hat{J}}^{b}]=i\,\epsilon_{abc}{\hat{J}}^{c}\;,$ (5.1)
where $\tau^{a}$ are the Pauli matrices, $a_{i},a^{\dagger}_{i}$ are the harmonic-oscillator annihilation and creation operators, of spin-up ($i=1$) and spin-down ($i=2$) particles,
$[a_{i},a^{\dagger}_{j}]=\delta_{ij}\;.$ (5.2)
[…]
The rotation matrix for a generic spin $j$ is simply (writing the Euler angles symbolically as $\Omega\equiv\{\gamma,\theta,\phi\}$) the direct productof the rotation matrices for spin $1/2$, (3.3),
$(D^{j}(\Omega))_{M,M^{\prime}}={\cal S}\left[(D^{1/2}(\Omega))_{m_{1},m_{1}^{\prime}}(D^{1/2}(\Omega))_{m_{2},m_{2}^{\prime}}\cdots(D^{1/2}(\Omega))_{m_{N},m_{N}^{\prime}}\right]$ (5.3)
symmetrized (${\cal S}$) with respect to $(m_{1},m_{2},\ldots,m_{N})$, where
$N=2j\;,\qquad M=\sum m_{i}\;,\qquad M^{\prime}=\sum m_{i}^{\prime}\;.$ (5.4)
It acts on the multiplet,
[…]
$|j,j-1\rangle$ $=$ $\frac{1}{\sqrt{N}}\{|\!\downarrow\rangle|\!\uparrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle+|\!\uparrow\rangle|\!\downarrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle+\ldots\}\;,$
$\ldots$
$|j,M\rangle$ $=$ $\frac{1}{\sqrt{{}_{N}C_{k}}}\{{\rm sum\,\,of\,\,terms\,\,with}\,\,k\,\,|\!\uparrow\rangle^{\prime}s\}\qquad(M=-j+k)$
$\ldots$
[…]
The action of $(D^{j}(\Omega))_{M,M^{\prime}}$ on particular states,
$|j,{\mathbf{n}}\rangle=|\tfrac{1}{2},{\mathbf{n}}\rangle\otimes|\tfrac{1}{2},{\mathbf{n}}\rangle\otimes\ldots|\tfrac{1}{2},{\mathbf{n}}\rangle\;,$ (5.6)
where $|\tfrac{1}{2},{\mathbf{n}}\rangle$ is given in Eq. (2.2), i.e. on the states with all spin $\tfrac{1}{2}$’s oriented in the same direction, is very simple. It is just the operation of rotating each and all spins towards the ${\hat{z}}$ direction (setting $\gamma=0$),
$(D^{j}(\Omega))|j,{\mathbf{n}}\rangle=|\!\uparrow\rangle|\!\uparrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle=(1,0,\ldots,0)^{T}=|j,{\hat{z}}\rangle\;.$ (5.7)
This corresponds precisely to the rotation of a vector,
$R(\Omega)\,{\bf j}^{(\rm class)}=R(\Omega)\,j\,{\mathbf{n}}=(0,0,j)^{T}\;,$ (5.8)
where $R(\Omega)$ is the standard $3\times 3$ rotation matrix of classical mechanics.
6 Conclusion
The observations made in this note render the concept that a spin ($S$) (or angular momentum, $j$) becomes classical in the limit, $S(j)\to\infty$, a more precise one. In particular, we found the identification between the states with minimum fluctuations ((3.11), (3.12), and (3.13)) with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of SG experiments, of the angular-momentum addition rule, and of the rotation matrix.
At the same time, our analysis has revealed a subtlety in the quantum-classical correspondence in this context. It reflects the difference between the spaces of the quantum-mechanical and classical angular momentum states of definite magnitude ($\mathbf{CP}^{2j}$ and $\mathbf{CP}^{1}\sim S^{2}$, respectively). It is the angular momentum states of minimum fluctuations, (3.11), (3.12), (3.13), that naturally smoothly transit into classical angular momentum in the $S(j)\to\infty$ limit, independently of temperatures. On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ are possible as pure quantum states only at exceedingly low temperatures. The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other. A macroscopic system is necessarily in a mixed state at finite temperatures [5, 6, 7],[1]. For instance, no macroscopically split pure wave packets in a SG setting survive decoherence effects.
The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles seems to be a fruitful one both mathematically and physically. From the mathematics point of view, the entire theory of angular momentum can be reconstructed this way [8], as we recalled already, and it helps to recover certain formulas easily. It looks also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized model of a macroscopic body, made of many atoms and molecules (spins). The sharp spike of the SG projection of the state $|S,{\bf n}\rangle$ in the narrow region around $S_{z}=S\cos\theta$ can then be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $S_{z}$.
The various form of suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is in fact analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [9], to explain the exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones.
Acknowledgments
The work by KK is supported by the INFN special-initiative project grant GAST (Gauge and String Theories). We are grateful to Hans-Thomas Elze, Jarah Evslin, Riccardo Guida, Pietro Menotti and Arkady Vainshtein for useful comments and precious advices.
Appendix A Newton’s equation for a macroscopic body
The conditions needed for the CM of an isolated macroscopic body at finite body temperatures to obey Newton’s equations have been discussed in great care by one of the present authors [1]. They are
  • (i)
    For macroscopic motions (for which $\hbar\simeq 0$) the Heisenberg relation does not limit the simultaneous determination – the initial condition – of the position and momentum;
  • (ii)
    The absence of quantum diffusion, due to a large mass (a large number of atoms and molecules composing the body);
  • (iii)
    A finite body temperature, implying the thermal decoherence and mixed-state nature of the body.
Under these conditions, the CM of an isolated macroscopic body has a unique trajectory. Newton’s equations for it follow from the Ehrenfest theorem. See Ref. [1] for various subtleties and for the explicit derivation of Newton’s equation under external gravitational forces, under weak, static, smoothly varying external electromagnetic fields, and under a harmonic-oscillator potential. Somewhat unexpectedly, the environment-induced decoherence [5, 6, 7] which is extremely effective in rendering macroscopic states in a finite-temperature environment a mixture, has been found not to be the most essential elements for the derivation of classical mechanics from quantum mechanics [1, 2].

### v3

@@
 Quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere.
-A spin $j$ (a $12j+$ 1-state system) for generic $j$ is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$.
-Here we study the angular momentum/spin in the limit, $j\to\infty$.
-The state, $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, where ${\hat{\mathbf{J}}}$ is the angular momentum operator and ${\mathbf{n}}$ stands for a generic unit vector in ${\bf R}^{3}$, is found to behave as a classical angular momentum, $j\,{\mathbf{n}}$.
-We discuss this phenomenon, by analysing the Stern-Gerlach experiments, the angular-momentum composition rule, and the rotation matrix.
-This problem arose from the consideration of a macroscopic body under an inhomogeneous magnetic field.
-Our observations help to explain how classical mechanics (with unique particle trajectories) emerges naturally from quantum mechanics in this context, and at the same time, make the widespread notion that large spins somehow become classical, a more precise one.
+A spin $j$ for a generic $j$ (a $12j+$ 1-state system) is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$.
+Here we study the state of a single angular momentum/spin in the limit, $j\to\infty$.
+The special class of states $|j,{\mathbf{n}}\rangle\in{\mathbf{CP}}^{2j}$, with spin oriented towards definite spatial directions ${\mathbf{n}}\in S^{2}$, i.e., $({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle$, are found to behave as classical angular momenta, $j\,{\mathbf{n}}$, in this limit.
+Vice versa, general spin states in ${\mathbf{CP}}^{2j}$ do not become classical, even at large $j$.
+We discuss these questions, by analysing the Stern-Gerlach processes, the angular-momentum composition rule, and the rotation matrix.
+Our observations help to clarify better how classical mechanics emerges from quantum mechanics in this context (e.g., with unique trajectories for a particle carrying a large spin), and to make the widespread idea that large spins somehow become classical, more precise.
 1 Introduction
-It is a widely shared view that in the limit of large spin (angular momentum) the quantum-mechanical spin somehow becomes classical.
-After all, the large spin limit ($S/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ limit, and, in accordance with the general idea of the semi-classical limit, physics of large angular momentum is expected to become that of classical angular momentum.
-But to the best of the authors’ knowledge, it has not been clearly exhibited in the literature how exactly this works  [^1] Understandably, there is a vast literature about large angular momentum, from the early days of quantum mechanics to today, in a wide range of research fields, in various contexts and problems, and also based on different motivations.
-In order to avoid being biased and necessarily incomplete, we shall refrain in this work from attempting to cite or comment on any of them..
-It is the aim of this note to fill in this gap.
-The question came under a sharp focus in a general discussion on emergence of classical mechanics for a macroscopic body from quantum mechanics [1, 2].
-Knowing about the typical Stern-Gerlach (SG) process for a spin $1/2$ atom, we ask what happens if a macroscopic body consisting of many (e.g.
+It is a widely shared view that in the limit of large spin (angular momentum) a quantum-mechanical spin somehow becomes classical [1]-[21].
+After all, the large spin limit ($j/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ ($j$  fixed) limit, and, in accordance with the general idea of the semi-classical limit in quantum mechanics, a large-angular-momentum state might be expected to behave as a classical angular momentum.
+But to the best of the authors’ knowledge, it has never been clearly shown whether this indeed occurs or if it does, exactly how.
+It is the aim of this note to help filling in this gap.
+The question was put under a sharp focus in a general discussion on emergence of classical mechanics for a macroscopic body from quantum mechanics, as formulated in [22, 23].
+Knowing about the typical Stern-Gerlach (SG) process for a spin $\frac{1}{2}$ atom, one asks what happens if a macroscopic body consisting of many (e.g.
 $N\sim\mathcal{O}(10^{23}))$ spins is set under an inhomogeneous magnetic field of a SG set-up.
-In general, with many spins oriented in different directions, the problem does not arise, as the bound-state Hamiltonian does not allow each component particle to split à la SG into spatially separated sub-wavepackets.
-Logically, however, one cannot exclude the special situation in which a large number of spins are all oriented in the same direction.
-Such a large spin system might continue to behave quantum-mechanically, e.g.
-with a typical SG spreading of the wavepackets, even in the $S/\hbar\to\infty$ limit.
-It will be shown that the state $|S,{\mathbf{n}}\rangle$, the state in which the large spin is oriented in some direction ${\mathbf{n}}$, behaves instead as a classical angular momentum, yielding the unique trajectory in the SG processes (Sec.
-2).
-This result leads us to consider more carefully in Sec.
-3 the relation between the space of quantum-mechanical angular-momentum states and that of classical angular momenta, and to identify the special class of quantum angular momentum states to be put in correspondence with the classical angular momentum (see Eqs.
-(3.11), (3.12), (3.13) below).
-The content of the present work is to check the consistency of this picture, through the analysis of the SG experiments (Sec.
-2), the angular-momentum composition rule (Sec.
-4), and the rotation matrix (Sec.
-5).
-We conclude with a brief reflection in Sec.
-6.
-2 Stern-Gerlach experiments
-2.1 Spin $1/2$ atom
-For definiteness let us take an incident atom with spin $S=\tfrac{1}{2}$ directed along a definite, but generic, direction ${\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, i.e.
-$\Psi=\psi({\bf r},t)\,|{\bf n}\rangle\;,$ (2.1)
-$|{\bf n}\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,\qquad c_{1}=e^{-i\phi/2}\cos\tfrac{\theta}{2}\;,\qquad c_{2}=e^{i\phi/2}\sin\tfrac{\theta}{2}\;.$ (2.2)
-When the atom passes through a region with an inhomogeneous magnetic field,
-${\mathbf{B}}=(0,0,B(z)),\qquad\frac{\partial B(z)}{\partial z}\neq 0\;,$ (2.3)
-the wave packet splits into two subpackets, with respective weights, $|c_{1}|^{2}$ and $|c_{2}|^{2}$: they manifest as the relative intensities of the image bands on the screen at the end.
-The original Stern-Gerlach experiment [3] has had a fundamental impact to the development of quantum theory, in proving the quantization of the angular momentum, the existence of half-integer spin, and more generally, demonstrating the quantum-mechanical nature of the center of mass (CM) of the silver atom [2].
-2.2 Large spin
 A classical particle with a magnetic moment directed towards
 ${\boldsymbol{\mu}}={\rm const}.\,{\bf n}\;,\qquad{\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;,$ (1.1)
-moves in an inhomogeneous magnetic field (2.3), according to Newton’s equations,
-$m\,{\dot{\bf r}}={\bf p}\;,\qquad\frac{dp_{x}}{dt}=\frac{dp_{y}}{dt}=0\;,\quad\frac{dp_{z}}{dt}=F_{z}=\frac{\partial}{\partial z}{\boldsymbol{\mu}}\cdot{\bf B}\;.$ (2.5)
-The way the unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [1], where the magnetic moment is the expectation value
+moves in an inhomogeneous magnetic field according to Newton’s equations,
+$m\,{\dot{\bf r}}={\bf p}\;,\qquad\frac{d{\bf p}}{dt}={\bf F}=\nabla({\boldsymbol{\mu}}\cdot{\bf B})\;.$ (1.2)
+The way how such a unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [22]. The magnetic moment of a macroscopic particle is the expectation value
 $\sum_{i}\langle\Psi|({\hat{\boldsymbol{\mu}}}_{i}+\frac{e_{i}{\hat{\boldsymbol{\ell}}}_{i}}{2m_{i}c})|\Psi\rangle={\boldsymbol{\mu}}\;,$ (1.3)
 taken in the internal bound-state wave function $|\Psi\rangle$ where $\mu_{i}$ and $\frac{e_{i}{\bf\ell}_{i}}{2m_{i}c}$ denote the intrinsic magnetic moment and the one due to the orbital motion of the $i$-th constituent atom or molecule ($i=1,2,\ldots,N$), respectively.
-Clearly, in general, the considerations made in Sec.
-2.1 for a spin $1/2$ atom, with a doubly split wavepacket, cannot be simply generalized to (or compared with) a classical body (2.6) with $N\sim\mathcal{O}(10^{23})$ constituents.
-Typically, the spins are oriented along different directions, but the particles carrying them are bound in the macroscopic body made of atomic, molecular and larger structures.
-The wave function of particles in bound states does not split in sub-wavepackets, as that of an isolated spin $\tfrac{1}{2}$ atom does, under an inhomogeneous magnetic field.
-The bound-state Hamiltonian does not allow it  [^2] This discussion is indeed closely parallel to the question of quantum diffusion.
+Clearly, in general, the well-known picture for a spin $\frac{1}{2}$ atom, with a doubly split wavepacket, does not generalize to a classical body (1.3) with $N\sim\mathcal{O}(10^{23})$ constituents.
+Typically, the spins are oriented along different directions, but the crucial point is that the component particles carrying them are bound in the macroscopic body in various atomic, molecular or crystalline structures.
+The wave function of particles in bound states does not split in sub-wavepackets, in contrast to that of an isolated spin $\tfrac{1}{2}$ atom, under an inhomogeneous magnetic field.
+The bound-state Hamiltonian does not allow it  [^1] This discussion is closely parallel to the question of quantum diffusion.
 Unlike free particles, particles in bound states (the electrons in atoms; atoms in molecules, etc.) do not diffuse, as they move in binding potentials.
 This is one of the elements for the emergence of the classical mechanics, with unique trajectories for macroscopic bodies.
-As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is suppressed by mass..
+As for the center-of-mass (CM) wave function of an isolated macroscopic body, its free quantum diffusion is simply suppressed by mass..
 What determines the motion of the CM of the body is the average magnetic moment, (1.3).
-Logically, nevertheless, one cannot exclude special cases (e.g.
-a magnetized piece of metal) of all spins of the body oriented in the same direction.
-One might wonder how the wave function of such a particle with spin $S\to\infty$ behaves, under an inhomogeneous magnetic field.
-The question is whether the conditions enlisted in Ref.
-[1] (see Appendix A), for the emergence of classical mechanics for a macroscopic body with a unique trajectory, are sufficient, or whether some extra condition (or a new mechanism) is required, to suppress possible wide spreading of the wave function into many sub-packets (see Fig.
-1 for a small spin, $S=13/2$) under an inhomogeneous magnetic field.
+Logically, however, one cannot exclude special cases (e.g.
+a magnetized piece of metal) with all spins inside the body oriented in the same direction.
+One might wonder how the wave function of such a particle with spin $j\to\infty$ behaves, under an inhomogeneous magnetic field.
+The question is whether the three conditions recognized in [22] for the emergence of classical mechanics for a macroscopic body with a unique trajectory, reviewed in Appendix A here, are indeed sufficient.
+Or, is some extra condition, or a new unknown mechanism, needed, to suppress possible wide spreading of the wave function into many sub-packets (such as in Fig.
+1 for a small spin, $j=13/2$) under an inhomogeneous magnetic field?
 Figure 1: The distribution $|c_{k}|^{2}$ in $k$, i.e.
-in possible values of $S_{z}$, $-S\leq S_{z}\leq S$, for a spin $S=13/2$ particle in the state (2.7), with $\theta=\pi/2$ (center, orange dots) and $\theta=\pi/4$ (right, blue dots).
-The answer turns out to be simple but somewhat unexpected.
-The following discussion, up to Eq.
-(2.15), is taken almost verbatim from Sec.
-3.1.3 of Ref.
-[2]; for the present work it represents the starting point, and serves as the main tool of analysis below.
-Consider the state of a spin $S$ directed towards a direction ${\mathbf{n}}$,
-$\left({\hat{\mathbf{S}}}\cdot{\mathbf{n}}\right)\,|S,{\mathbf{n}}\rangle=S\,|S,{\mathbf{n}}\rangle\;,\qquad{\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (2.7)
-Choosing the magnetic field (and its gradients) in the ${z}$-direction, we need to express $|S,{\mathbf{n}}\rangle$ as a superposition of the eigenstates of ${\hat{S_{z}}}$ ($S=N/2$),
-$|S,{\mathbf{n}}\rangle=\sum_{k=0}^{N}c_{k}\,|S,S_{z}=M\rangle\;,\qquad M=-\frac{N}{2}+k\;,\quad(k=0,1,\ldots,N)\;.$ (2.8)
-The expansion coefficients $c_{k}$ are known:
-$c_{k}={\binom{N}{k}}^{1/2}\,e^{-iM\phi/2}\left(\cos\tfrac{\theta}{2}\right)^{k}\left(\sin\tfrac{\theta}{2}\right)^{N-k}\;,\qquad\sum_{k=0}^{N}|c_{k}|^{2}=1\;,$ (2.9)
-where $\binom{N}{k}=N!/k!(N-k)!$ are the binomial coefficients.
-To get (2.9), let us consider (2.7) as a direct product of $N$ spin $\tfrac{1}{2}$ particles, (2.2),
-$|S,{\mathbf{n}}\rangle=|{\bf n}\rangle_{1}\otimes|{\bf n}\rangle_{2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (2.10)
-all oriented in the same direction. Expanding and collecting terms with a fixed $k$ (the number of spin-up particles) one arrives at Eq. (2.9).
-Using Stirling’s formula, one can derive, for large values of both $N$ and $k$ with $x=k/N$ fixed, the distribution across different values of $S_{z}=M$,
-$|c_{k}|^{2}\simeq e^{Nf(x)}\;,\qquad x=k/N\;,$ (2.11)
-with
-$f(x)=-x\log x-(1-x)\log(1-x)+2x\log\cos\tfrac{\theta}{2}+2(1-x)\log\sin\tfrac{\theta}{2}\;.$ (2.12)
-The saddle-point approximation, valid at $N\to\infty$, yields
-$f(x)\simeq-\frac{(x-x_{0})^{2}}{x_{0}(1-x_{0})}\;,\qquad x_{0}=\cos^{2}\tfrac{\theta}{2}\;,$ (2.13)
-and thus
+in possible values of $J_{z}=m$, $-j\leq m\leq j$ ($m=-j+k$) for a spin $j=13/2$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange dots) or with $\theta=\pi/4$ (right, blue dots).
+The aim of this work is to answer these questions.
+We shall first discuss in Sec.
+2 some mathematical and physical aspects of the quantum states of a small and large angular momenta, and compare them with the properties of a classical angular momentum.
+Considerations of quantum fluctuations of various angular momentum states lead us to propose that a particular class of states, $|j,{\mathbf{n}}\rangle$,
+${\hat{\mathbf{J}}}^{2}\,|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\,,\qquad({\hat{\mathbf{J}}}\cdot{\mathbf{n}})\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\,,$ (1.4)
+where ${\mathbf{n}}$ is a unit vector directed towards $(\theta,\phi)$ direction,
+${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (1.5)
+are to be identified with a classical angular momentum, $j\,{\mathbf{n}}$, in the large $j$ limit.
+The states (1.4) are known in the literature as coherent spin states or Bloch states [5, 10, 6, 11, 7].
+Vice versa, generic states of large spin will be found to remain quantum mechanical, with large fluctuations, even in the limit, $j\to\infty$, showing that the often stated belief that a large spin (angular momentum) becomes classical, should not be taken for granted, literally.
+In subsequent sections, we are going to examine these ideas through the analysis of the Stern-Gerlach processes (Sec.
+3), the orbital angular momenta (Sec.
+4), the angular-momentum addition rule (Sec.
+5), and the rotation matrix (Sec.
+6).
+Generic spin $j$ states far from (1.4) are shown to remain quantum mechanical even in the limit $j\to\infty$: the fate of these states will be discussed in Sec.
+7.
+Conclusion and a few more general reflections are in Sec.
+8.
+2 Space of quantum spin states and classical angular momenta
+The generic pure spin $\frac{1}{2}$ (two-state system) state is described by the wave function
+$|\psi\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,$ (2.1)
+where
+$(c_{1},c_{2})\sim\lambda(c_{1},c_{2})\;,\qquad\lambda\in{\mathbf{C}}/\{0\}\,$ (2.2)
 […]
-in the $N\to\infty$ ($x=k/N$ fixed) limit. The narrow peak position $x=x_{0}$ corresponds to (see Eq. (2.8))
-$S_{z}=M=N(x-\tfrac{1}{2})=S\,(2\cos^{2}\tfrac{\theta}{2}-1)=S\,\cos\theta\;.$ (2.15)
-This means that a large spin ($S\gg\hbar$) quantum particle with spin directed towards ${\mathbf{n}}$, in a Stern-Gerlach setting with an inhomogeneous magnetic field (2.3), moves along a single trajectory of a classical particle with $S_{z}=S\,\cos\theta$, instead of spreading over a wide range of split sub-packet trajectories covering $-S\leq S_{z}\leq S$.
-See Fig.
-2, cfr.
-Fig.
-1.
-Figure 2: The distribution $|c_{k}|^{2}$ in $k$, i.e.
-in possible values of $S_{z}$, for a spin $S=1\cdot 10^{3}$ particle in the state (2.7), with $\theta=\pi/2$ (center, orange peak) and $\theta=\pi/4$ (right, blue peak).
-The particle starts looking like classical.
-This (perhaps) somewhat surprising result shows that quantum mechanics (QM) takes care of itself, so to speak, in ensuring that a large spin particle ($S/\hbar\to\infty$) traces a classical trajectory.
-This is consistent with the known general behavior of the wave function in the semi-classical limit ($\hbar\to 0$)  [^3] Of course this does not mean that the classical limit necessarily requires or implies $S/\hbar\to\infty$, though..
-Note that if the value $S\gg\hbar$ is understood as due to the large number of spin $1/2$ particles composing it (see the comment after (2.9)), the spike (2.13), (2.14) can be understood as due to the accumulation of a huge number of microstates giving $S_{z}\simeq M$.
-Another, equivalent way to see the shrinkage of the distribution in $S_{z}$, (2.14), is to study its dispersion.
-Given the relative weights $|c_{k}|^{2}$, (2.9), where $S_{z}=-S+k$ and $S=\tfrac{N}{2}$, with $k=0,1,\ldots,N$, it is a simple exercise to find that
-$\langle S_{z}\rangle=S\,\cos\theta\;,\qquad\langle S_{z}^{2}\rangle=S^{2}\cos^{2}\theta+\frac{S}{2}\sin^{2}\theta\;,$ (2.16)
-so that the dispersion (the fluctuation width) is given by
-$(\Delta S_{z})^{2}=\langle(S_{z}-\langle S_{z}\rangle)^{2}\rangle=\langle S_{z}^{2}\rangle-\langle S_{z}\rangle^{2}=\frac{S}{2}\sin^{2}\theta\;.$ (2.17)
-It follows that
-$\lim_{S/\hbar\to\infty}\frac{|\Delta S_{z}|}{S}=0\;,$ (2.18)
-i.e.
-it is an infinitely narrow Gaussian.
-Note that the SG set-up is indeed a measurement of $S_{z}$ in the state (2.7).
-The fact that the result is always $S_{z}\sim\langle S_{z}\rangle=S\,\cos\theta$ with no dispersion, means that it is a classical angular momentum $\sim S{\mathbf{n}}$ (see Eqs.
-(3.11), (3.12) below).
-2.2.1 Density matrix
-The same phenomenon can be seen also by using the density matrix. Being the state (2.7) a pure state, its $(2S+1)\times(2S+1)$ density matrix is given by
-$\rho_{k\ell}=c_{k}c_{\ell}^{*}\,,$ (2.19)
-where $c_{k}(c_{\ell})$ are given in Eq.
-(2.9).
-In the limit as $S\to\infty$, it can be observed that it essentially reduces to a single diagonal element, provided that the fluctuations of the order of $\sqrt{S}$ are neglected in comparison to $S$.
-The density matrix ${\boldsymbol{\rho}}$ for $S=500$, with $(\theta,\phi)=(\frac{\pi}{2},0)$ and $(\theta,\phi)=(\frac{\pi}{4},0)$, is depicted in Figs.
-3.
-Figure 3: The density matrix ${\boldsymbol{\rho}}$ for the state (2.7), of spin $S=5\cdot 10^{2}$, with ($a$): $(\theta,\phi)=(\tfrac{\pi}{2},0)$ and ($b$): $(\theta,\phi)=(\tfrac{\pi}{4},0)$.
-The height of the plots indicates the matrix elements $\rho_{k\ell}$.
-Note that ${\boldsymbol{\rho}}$ is real for $\phi=0$.
-2.2.2 The state preparation and successive SG set-ups
-For a small $S$, as it is well known, the SG set-up can be used to prepare any desired state $|S,S_{z}=M\rangle$, by an intentionally biased SG apparatus, e.g.
-in which all the split sub-wavepackets are blocked, except for the one corresponding to $|S,S_{z}=M\rangle$.
-Equivalently, one may use the so-called null-measurements (see Ref.
-[4] for a recent review).
-In any case, for a small finite $S$, where there are $12S+$ 1 well-separated sub-packets (see Fig.
-1), there is no difficulty in extracting any desired quantum state for subsequent experiments (state preparation).
-In the large-spin limit ($S\to\infty$) we are considering, since there is essentially only one wavepacket present (see (2.15)), there is no possibility of extracting or selecting the desired sub-wavepacket.
-The experimental resolution would not allow it.
-We have in mind a macroscopic scenario where, in comparison to the finite region (in principle) occupied by the waves $-S\leq M\leq S$, the fluctuations of the values of $M$ around (2.15) are of the order of $\sqrt{S}\ll S$, i.e.
-of zero width.
-We might consider introducing a second set of Stern-Gerlach (SG) magnetic fields directed in a different direction, denoted by ${\hat{z}}^{\prime}$, as a new quantization axis.
-This setup allows us to study the state emerging from the first set of SG magnetic fields (without the presence of an imaging screen),
-$|S,{\mathbf{n}}\rangle\sim c_{k}\,|S,S_{z}=M\rangle\;,\qquad k\sim 2S\cos^{2}\tfrac{\theta}{2}\;,\qquad M\sim S\,\cos\theta\;.$ (2.20)
-Since the sum of states around $k\sim 2S\cos^{2}\left(\frac{\theta}{2}\right)$ is essentially equal to the complete sum, this state can be rewritten in terms of the eigenvalues $M^{\prime}=-S+k^{\prime}$, where $k^{\prime}=0,1,\ldots 2S$ correspond to the eigenvalues of $S_{z^{\prime}}$ as:
-$|S,{\mathbf{n}}\rangle$ $=$ $\sum_{k,k^{\prime}}c_{k}\,|S,M^{\prime}\rangle\langle S,M^{\prime}|S,M\rangle$ (2.21)
-$=$ $\sum_{k,k^{\prime}}|S,M^{\prime}\rangle\langle S,M^{\prime}|S,M\rangle\langle S,M|S,{\mathbf{n}}\rangle$
-$=$ $\sum_{k^{\prime}}|S,M^{\prime}\rangle\langle S,M^{\prime}|S,{\mathbf{n}}\rangle=\sum_{k^{\prime}}c_{k^{\prime}}|S,M^{\prime}\rangle\;.$
-This represents an expansion of the original state $|S,{\mathbf{n}}\rangle$ as a linear combination of the eigenstates of $S_{z^{\prime}}$.
-In the limit $S\to\infty$, the results from Section 2.2 tell us that the sum is dominated by the small region around $S_{z^{\prime}}\sim S,\cos({\mathbf{n}}\cdot{\hat{z}}^{\prime})$, with negligible fluctuations.
-The trajectory remains unique, regardless of the choice of quantization axis (i.e.
-the direction of the magnetic field) ${\hat{z}}^{\prime}$.
-2.2.3 States almost directed towards ${\vec{\mathbf{n}}}$
-Obviously, in the large spin limit, other states close to $|S,{\mathbf{n}}\rangle$,
-$\left({\hat{\mathbf{S}}}\cdot{\mathbf{n}}\right)\,|S,{\mathbf{n}},q\rangle=(S-q)\,|S,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,\,\,{\rm fixed}$ (2.22)
-(i.e. the spin roughly oriented towards ${\mathbf{n}}$) will behave similarly. For instance, the $q=1$ state is (cfr. Eq. (2.10))
-$|S,{\mathbf{n}},1\rangle$ $=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}\,|{\bf n}\rangle_{1}\otimes\cdots|{\bf n}\rangle_{\ell}\otimes|-{\bf n}\rangle\otimes|{\bf n}\rangle_{\ell+2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (2.23)
-$=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}|S_{1},{\mathbf{n}}\rangle\otimes|-{\bf n}\rangle\otimes|S_{2},{\mathbf{n}}\rangle\;,$
-where $|S_{1},{\mathbf{n}}\rangle$ and $|S_{2},{\mathbf{n}}\rangle$ ($S_{1}=\tfrac{\ell}{2}\;,S_{2}=\tfrac{N-\ell-1}{2}$), are two spins oriented towards ${\mathbf{n}}$, (2.8), as already studied in Sec.
-2.2.
-In the $S\to\infty$ limit, the sum is dominated by terms with $\ell\sim\mathcal{O}(N)\to\infty$.
-The SG projection of this state onto the $S_{z}$ eigenstates is therefore approximated by
-$S_{z}\sim S_{1\,z}+S_{2\,z}=S_{1}\cos\theta+S_{2}\cos\theta\simeq S\,\cos\theta\;,$ (2.24)
-where the results of Sec.
-2.2 have been used.
-A similar result holds for $|S,{\mathbf{n}},q\rangle$ for any finite $q$, showing that all these states behave as a classical angular momentum, $S\,{\mathbf{n}}$.
-3 Quantum and classical angular momentum states
-Before proceeding, we must discuss the correspondence between the quantum mechanical angular momentum states and their classical counterparts, more carefully.
-The spin $1/2$ (two-state system) wave function is given by
-$|\psi\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,$ (3.1)
-where $(c_{1},c_{2})\sim\lambda(c_{1},c_{2})$ with $\lambda\in{\mathbf{C}}/\{0\}$, describes the points of $\mathbf{CP}^{1}\sim S^{2}$.
-We note that for a spin $1/2$, any pure state (3.1) can be interpreted as the state in which the spin is oriented in some space direction ${\mathbf{n}}$, as in Eq.
-(2.2), without loss of generality.
-This is because the manifold of pure spin $1/2$ quantum states and that of the unit space vector are both $S^{2}\sim\mathbf{CP}^{1}$.
+${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;,$ (2.3)
+that is
+${\hat{J}}^{2}\left|\frac{1}{2},{\mathbf{n}}\right\rangle=\frac{3}{4}\left|\frac{1}{2},{\mathbf{n}}\right\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,\left|\frac{1}{2},{\mathbf{n}}\right\rangle=\frac{1}{2}\,\left|\frac{1}{2},{\mathbf{n}}\right\rangle\;,$ (2.4)
+$|{\bf n}\rangle\equiv\left|\frac{1}{2},{\mathbf{n}}\right\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,\qquad c_{1}=e^{-i\phi/2}\cos\tfrac{\theta}{2}\;,\qquad c_{2}=e^{i\phi/2}\sin\tfrac{\theta}{2}\;.$ (2.5)
+without loss of generality.
+This is so because the space of pure spin $\frac{1}{2}$ quantum states and that of the unit space vector coincide: they are both $S^{2}\sim\mathbf{CP}^{1}$.
 Indeed, given the state (2.5), there is always a rotation matrix such that
-$D^{1/2}(0,\theta,\phi)\cdot\left(\begin{array}[]{c}e^{-i\phi/2}\cos\tfrac{\theta}{2}\\ e^{i\phi/2}\sin\tfrac{\theta}{2}\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\;.$ (3.2)
-We recall that the general rotation matrix for a spin $1/2$ is given, using Euler angles  [^4] Note that the third Euler angle (the rotation angle $\gamma$ about the final $z$-axis) is redundant here, as it gives only the phase $e^{i\gamma/2}$, and has been set to 00 in (3.2).
-($\gamma,\theta,\phi$), as follows
+$D^{1/2}(0,\theta,\phi)\cdot\left(\begin{array}[]{c}e^{-i\phi/2}\cos\tfrac{\theta}{2}\\ e^{i\phi/2}\sin\tfrac{\theta}{2}\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\;,$ (2.6)
+where the general rotation matrix for a spin $\frac{1}{2}$ is given, using Euler angles  [^2] Note that the third Euler angle (the rotation angle $\gamma$ about the final $z$-axis) is redundant here, as it gives only the phase $e^{i\gamma/2}$, and has been set to 00 in Eq.(2.6).
+($\gamma,\theta,\phi$), by
 $D^{1/2}(\gamma,\theta,\phi)=U_{z}(\gamma)U_{y}(\theta)U_{z}(\phi)=\left(\begin{array}[]{cc}e^{i(\gamma+\phi)/2}\cos\tfrac{\theta}{2}&e^{(\gamma-\phi)/2}\sin\tfrac{\theta}{2}\\ -e^{i(-\gamma+\phi)/2}\sin\tfrac{\theta}{2}&e^{-i(\gamma+\phi)/2}\cos\tfrac{\theta}{2}\end{array}\right)\;.$ (2.7)
-The situation is different for a generic spin $j=N/2$  [^5] In this section we denote the general angular momentum (including orbital angular momentum) by $j$, rather than $S$, a symbol adopted commonly to indicate the spin (the intrinsic angular momentum of a particle).
-In this work we will use both $j$ and $S$.
-Likewise, we shall use the terms “spin” and “angular momentum”, interchangeably..
-The wave function now has the form,
-$|\psi\rangle=\sum_{k=0}^{N}b_{k}|j,j_{z}=-j+k\rangle$ (3.4)
+The fact that any spin $\frac{1}{2}$ quantum state can be associated with a definite space direction, however, does not mean that it can be identified with a classical angular momentum, $\frac{1}{2}\hbar{\mathbf{n}}$, as the fluctuations in the directions transverse to ${\mathbf{n}}$ are always of the same order of the spin magnitude itself (see Eq.(2.13) below, with $j=\frac{1}{2}$, $m=\pm\frac{1}{2}$).
+It is always a quantum mechanical system.
+Let us now consider a generic spin $j$ state. It is a special type of $12j+$ 1-state quantum system. Its general wave function has the form,
+$|\psi\rangle=\sum_{m=-j}^{j}b_{m}|j,j_{z}=m\rangle$ (2.8)
 where
-$(b_{0},b_{1},\ldots,b_{N})\sim\lambda(b_{0},b_{1},\ldots,b_{N})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (3.5)
-are the coordinates of the space $\mathbf{CP}^{N}$.
+$(b_{-j},b_{-j+1},\ldots,b_{j})\sim\lambda(b_{-j},b_{-j+1},\ldots,b_{j})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (2.9)
+are the coordinates of the complex projective space $\mathbf{CP}^{2j}$ 333 As is well known, the manifold of pure quantum states of any $n$-state system is $\mathbf{CP}^{n-1}$ [24].
+The spin states (2.9) are special cases with $n=2j+1$.
+.
 On the other hand, the variety of the directions in the three-dimensional space, ${\mathbf{n}}$, is always $S^{2}\sim\mathbf{CP}^{1}$.
-It means that the state of the spin oriented towards a definite direction (2.8), is not a general pure spin $j$ state, but belongs to a small subset, $\mathbf{CP}^{1}\subset\mathbf{CP}^{2j}$.
-In other words, not all states of spin $j$, as described in (3.4), can be transformed by a rotation matrix (selecting an appropriate new ${\hat{z}}$ axis) into the form:
-$(1,0,\ldots,0)^{T}\;.$ (3.6)
-These considerations highlight a potential subtlety in our argument that quantum mechanical angular momentum states somehow become classical in the large $j$ limit.
-Let us recall that because of the commutation relations
+It means that this time not all states of spin $j$ state, (2.8), can be transformed by a rotation matrix (selecting an appropriate new ${\hat{z}}$ axis) into the form:
+$(b_{-j}^{\prime},b_{-j+1}^{\prime},\ldots,b_{j}^{\prime})\sim(0,0,\ldots,0,1)\;.$ (2.10)
+This observation is the first hint that there are some subtleties in the way classical picture of angular momentum emerges from quantum mechanics at large $j$.
+It is the purpose of the present note to elaborate on this point.
+Let us start with the basic properties of quantum mechanical angular momentum. These are of course well-known textbook materials. Because of the commutation relations
 $[{\hat{J}}_{k},{\hat{J}}_{\ell}]=i\hbar\,\epsilon^{k\ell m}{\hat{J}}_{m}\;,$ (2.11)
@@
 In the state, $|j,m\rangle$, where ${\hat{J}}_{z}$ has a definite value, $m\,\hbar$, ${\hat{J}}_{x}$ and ${\hat{J}}_{y}$ are fluctuating.
-Their magnitudes $\Delta{J}_{x,y}=\langle j,m|{\hat{J}}_{x,y}^{2}|j,m\rangle^{1/2}$ are found to be
+Their magnitudes $\Delta{J}_{x,y}=\langle j,m|{\hat{J}}_{x,y}^{2}|j,m\rangle^{1/2}$ are given by
 $\Delta{J}_{x}=\Delta{J}_{y}=\left(\langle j,m|\big{(}{\hat{J}}^{2}-{\hat{J}}_{z}^{2}\big{)}|j,m\rangle/2\right)^{1/2}=\left(j(j+1)-m^{2})/2\right)^{1/2}\hbar\sim\mathcal{O}(j)\hbar\;.$ (2.13)
-Therefore, in a generic state $|j,m\rangle$, there are strong quantum fluctuations of $J_{x}$ and $J_{y}$ at $j\to\infty$.
+Namely, in a generic state $|j,m\rangle$, there are strong quantum fluctuations of $J_{x}$ and $J_{y}$ at $j\to\infty$, of the same order of magnitude with $j$ itself.
 There is no way such a state can be associated with a classical angular momentum, which requires all three components to be well-defined simultaneously.
@@
 $\Delta{\hat{J}}_{x}=\Delta{\hat{J}}_{y}=(j/2)^{1/2}\hbar\ll j\,\hbar\;.$ (2.14)
-For these state, it makes sense to interpret them “classically” as angular momentum pointed towards ±$\pm{\hat{z}}$, as its fluctuations $(\sim\mathcal{O}(\sqrt{j})$) become negligible with respect to its magnitude $j$, in the limit $j\to\infty$.
-The same holds true in all states in which spin is oriented towards a generic direction ${\mathbf{n}}$, as has been already noted in the previous section, Eqs.
-(2.16)-(2.18).
-Accordingly, we propose the correspondence (by setting $\hbar=1$)
-$\boxed{\quad|j,{\mathbf{n}}\rangle\quad\longleftrightarrow\quad{\bf j}^{(\rm class)}=j\,{\bf n}\;\quad}$ (3.11)
+For these state, it makes sense to interpret them as a classical angular momentum pointed towards ±$\pm{\hat{z}}$, as its transverse fluctuations $(\sim\mathcal{O}(\sqrt{j})$) are negligible with respect to its magnitude $j$, in the limit $j\to\infty$.
+The same can be said of the states of spin almost oriented towards $\hat{z}$, $|j,\pm(j-r)\rangle$, $r\ll j$.
+Naturally, all states of the form, $|j,{\mathbf{n}}\rangle$, in which spin is oriented towards a generic direction ${\mathbf{n}}$,
+${\hat{J}}^{2}|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\;,$ (2.15)
+have this property: its fluctuations in the components perpendicular to ${\mathbf{n}}$ become negligible in the $j\to\infty$ limit.
+The states (2.15) are known as Bloch states or coherent spin states in the literature [5, 10, 6, 11, 7].
+These observations lead us to propose that the quantum-classical correspondence (by setting $\hbar=1$) to be made is
+$|j,{\mathbf{n}}\rangle\quad\longleftrightarrow\quad{\bf j}^{(\rm class)}=j\,{\bf n}\;.$ (2.16)
+Of course, all states “almost oriented towards ${\mathbf{n}}$”, $|j,{\mathbf{n}},q\rangle,$
+$\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,,\,\quad q\ll j$ (2.17)
+discussed in Sec.
+3.2.2 (see Eq.(3.17)) below, can also be interpreted as ${\bf j}^{(\rm class)}=j\,{\bf n}$.
+We are going to examine below the consistency of such a picture, with respect to the SG processes (Sec.
+3 below), a study of orbital angular momentum (Sec.
+4), the addition rule of the angular momenta (Sec.
+5) and the rotation matrix (Sec.
+6).
+A powerful idea, useful both for mathematical and physical reasoning below, is that the general spin $j$ can be regarded as a direct product state of $n=2j$ spins, each carrying 2Planck-constant-over-$\frac{\hbar}{2}$.
+From a group-theoretic point of view, it is well known that any irreducible representation of the SU(2) group, which is the double covering group (see e.g., [25]) of the rotation group SO(3), is associated with the Young tableaux,
+[…]
+with a single row of $n$ boxes, i.e., the totally symmetric direct product of $n=2j$ spin $\frac{1}{2}$ objects.
+Indeed, the entire theory of quantum angular momentum can be reconstructed based on this point of view [26], by using the operators,
+${\hat{J}}^{a}=\sum_{i,j=1,2}a_{i}^{\dagger}\Big{(}\frac{\tau^{a}}{2}\Big{)}_{ij}a_{j}\;,\qquad[{\hat{J}}^{a},{\hat{J}}^{b}]=i\,\epsilon_{abc}{\hat{J}}^{c}\;,$ (2.19)
+where $\tau^{a}$ are the Pauli matrices, $a_{i},a^{\dagger}_{i}$ are the harmonic-oscillator annihilation and creation operators, of spin-up ($i=1$) and spin-down ($i=2$) particles,
+$[a_{i},a^{\dagger}_{j}]=\delta_{ij}\;.$ (2.20)
+[…]
+Physically, on the other hand, one may consider (2.18) as a particular state of a toy-model macroscopic body made of $n$ atoms, each carrying spin $\frac{1}{2}$.
+Now a particularly attractive picture of the quantum states (2.18) of general spin $j$, follows from the so-called stellar representation of $\mathbf{CP}^{n}$ ($n=2j$) [24]. Let the complex numbers
+$(Z_{0},Z_{1},\ldots,Z_{n})\sim\lambda\,(Z_{0},Z_{1},\ldots,Z_{n})\;,\qquad\lambda\in{\mathbf{C}}/{\{0\}}$ (2.21)
+be the homogeneous coordinates of $\mathbf{CP}^{n}$, and consider the $n$-th order equation,
+$\sum_{k=0}^{n}Z_{k}\,z^{n-k}=Z_{0}\,\prod_{\ell=1}^{n}(z-w_{\ell})=0\;.$ (2.22)
+As the equation (2.22) is invariant under the rescaling (2.21), the collection of the $n$ roots for $z$, $\{w_{\ell}\}$, can each be regarded as equivalent to a point of $\mathbf{CP}^{n}$.
+Note that no generality is lost by considering the neighborhood $Z_{0}\neq 0$ above: for instance, one may introduce the variable $z^{\prime}\equiv 1/z$ and rewrite equation (2.22) in terms of $z^{\prime}$.
+One sees then that the equation with $Z_{0}=0$ contains (at least one) root $z^{\prime}=0$, that is $z=\infty$.
+The explicit relation between $\{w_{\ell}\}$ and the $\mathbf{CP}^{n}$ point can be found by working backward: given any collection of $n$ roots $\{w_{\ell}\}$ in the complex plane, one can find the coefficients $\{Z_{k}\}$ the coordinates of $\mathbf{CP}^{n}$ by simply expanding the second expression of (2.22).
+The connection between this construction and the picture (2.18) above comes from the fact that any complex number $\{w_{\ell}\}$ can be regarded as the stereographic projection of a point on a sphere $S^{2}$, from the north (or south) pole, onto the plane containing the equator.
+See Fig.
+2.
+Naturally, a point on $S^{2}$ can be identified with the unit vector ${\mathbf{n}}$, (2.3), the orientation of each component spin $\frac{1}{2}$.
+The explicit relation is
+$w_{\mathbf{n}}=\frac{e^{i\phi}}{\tan(\theta/2)}\;.$ (2.23)
+Collecting things, we see that a point in $\mathbf{CP}^{n}$ is represented by an unordered collection of $n$ points on a sphere (stars), see Fig.
+3.
+This is known as the stellar representation of $\mathbf{CP}^{n}$.
+Figure 2: Two points (stars) on the sphere, and their stereographic projections from the north pole (N.P.) onto the plane containing the equator.
+Figure 3: A general point in $\mathbf{CP}^{n}$ is equivalent to an unordered collection of $n$ points (stars) on a sphere, in the stellar representation, (2.22).
+Each star is associated with the complex number $w_{\ell}$ by a stereographic projection, as in Fig.
+2.
+From the group-theoretic point of view, (2.18), a generic quantum state of spin $j$, is equivalent to the collection (the direct product) of $n=2j$ spin $\frac{1}{2}s,$
+${\cal S}\,\,|{\bf n}_{1}\rangle|{\bf n}_{2}\rangle\ldots|{\bf n}_{n}\rangle$ (2.24)
+where ${\cal S}$ stands for a symmetrization, orientated towards various directions ${\mathbf{n}}_{k}$, $k=1,2,\ldots,n$, see Eq.(2.5).
+(2.24) represents a general state of spin $j=2n$ and naturally corresponds to the random sets of stars in Fig.
+3.
+The particular state in which the spin $j$ is oriented towards a definite direction ${\mathbf{n}}$, (2.15), corresponds to the case in which the component $\frac{1}{2}$ spins are all oriented towards the same direction ${\mathbf{n}}_{1}={\mathbf{n}}_{2}=\ldots={\mathbf{n}}_{n}={\mathbf{n}}$,
+$|j,{\mathbf{n}}\rangle=|{\bf n}\rangle|{\bf n}\rangle\ldots|{\bf n}\rangle$ (2.25)
+(which is automatically symmetric).
+See Fig.
+4.
+By using (2.5) for each $|{\bf n}\rangle$, and collecting terms with the same fixed number $k$ of the spin up factors, we find $|j,{\mathbf{n}}\rangle$ as an expansion in terms of the eigenstates of $J_{z}$ with $J_{z}=m$,
+$|j,{\mathbf{n}}\rangle=\sum_{k=0}^{2j}c_{k}\,|j,m\rangle\;,\qquad m=-j+k\;,$ (2.26)
+$c_{k}={\binom{2j}{k}}^{1/2}\,e^{i(j-k)\phi}\left(\cos\tfrac{\theta}{2}\right)^{k}\left(\sin\tfrac{\theta}{2}\right)^{2j-k}\;,\qquad\sum_{k=0}^{2j}|c_{k}|^{2}=1\;,$ (2.27)
+where $\binom{n}{k}=n!/k!(n-k)!$ are the binomial coefficients.
+The states (2.26), (2.27), coincide, apart from an overall phase factor, with the Bloch (or coherent spin) states defined in [5, 10, 6, 11, 7], and denoted as $|j,\Omega\rangle$, $|j,\{\theta,\phi\}\rangle$, etc.
+In order to understand this formula from the point of view of the stellar representation of $\mathbf{CP}^{n}$, we need to translate the homogeneous coordinates of $\mathbf{CP}^{n}$ to the coefficients $\{c_{k}\}$ in (2.26) as
+$Z_{k}=c_{k}^{2}\;,\qquad k=0,1,\ldots,n\;,$ (2.28)
+and at the same time replace the projection of the stars on $S^{2}$ as
+$w_{\ell}\to w_{\ell}^{2}\;,\qquad\forall\ell\;.$ (2.29)
+Eq.(2.28) and (2.29) reflect the double covering of SO(3) group by SU(2).
+In the stellar representation of $\mathbf{CP}^{n}$, (2.21), (2.22), the state (2.25) corresponds to the situation where all roots coincide,
+$w_{\ell}=w_{\mathbf{n}}\;,\qquad\forall\ell\,.$ (2.30)
+All stars are at a coincident point ${\mathbf{n}}$ on the sphere, see Fig. 4.
+Figure 4: The coherent spin state (2.15), (2.26), (2.27) in the stellar representation.
+Taking into account the doubling (2.28) and (2.29) and expanding ($n=2j$)
+$\sum_{k=0}^{n}(c_{k})^{2}\,z^{n-k}=\prod_{\ell=1}^{n}(z-w_{\ell}^{2})=(z-w_{\mathbf{n}}^{2})^{n}\;,$ (2.31)
+($w_{\mathbf{n}}=\frac{e^{i\phi}}{\tan(\theta/2)}$), one finds precisely (2.26), (2.27), apart from unobservable phases and an overall factor.
+Thus the stellar representation of $\mathbf{CP}^{n}$ points - the generic states of spin $j$ provides us with an appealing intuitive picture of quantum fluctuations.
+The $\mathbf{CP}^{n}$ points with randomly distributed stars, Fig.
+3, represent spin states with large quantum fluctuations.
+The special states, (2.15), (2.25) - (2.27) with coincident stars, Fig.
+4, instead, are those with the minimum fluctuations, and correspond to the Bloch or spin coherent state [5, 10, 6, 11, 7].
+In the next four sections we shall illustrate how these latter states effectively behave as the classical angular momenta, $j\,{\mathbf{n}}$, in the $j\to\infty$ limit.
+3 Stern-Gerlach experiments
+In this section we compare the Stern-Gerlach experiment [27] for small and large spins, with special attention to the Bloch states, (2.15), (2.26), (2.27).
+3.1 Spin $1/2$ atom
+For definiteness let us take an incident atom with spin $j=\tfrac{1}{2}$ directed along a definite, but generic, direction ${\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, i.e.
+$\Psi=\psi({\bf r},t)\,|{\bf n}\rangle\;,$ (3.1)
+$|{\bf n}\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,\qquad c_{1}=e^{-i\phi/2}\cos\tfrac{\theta}{2}\;,\qquad c_{2}=e^{i\phi/2}\sin\tfrac{\theta}{2}\;.$ (3.2)
+Let us suppose that the atom enters from $x=-\infty$ and moves towards $x=+\infty$. When the atom passes through a region with an inhomogeneous magnetic field with
+$\frac{\partial B_{z}}{\partial z}\neq 0\;,$ (3.3)
+the wave packet splits into two subpackets, with respective weights, $|c_{1}|^{2}$ and $|c_{2}|^{2}$: they manifest as the relative intensities of the two separate image bands on the screen at the end.
+[…]
+The original Stern-Gerlach experiment (1922) [27] has had a fundamental impact to the development of quantum theory, in proving the quantization of the angular momentum, the existence of half-integer spin, and more generally, demonstrating the quantum-mechanical nature of the silver atom as a whole [23].
+The wave function of the form (3.2) corresponds to a pure state, i.e., $100$% polarized beam where all incident atoms are in the same spin state.
+If the beam is partially polarized or unpolarized, the state is described by a generic density matrix $\rho$, with
+$\mathop{Tr}\nolimits\,\rho=1\;,\qquad\rho_{ii}\geq 0\;,\qquad i=1,2\;.$ (3.4)
+The Stern-Gerlach experiment measures the relative frequency that the atom arriving at the screen happens to have spin $s_{z}=\tfrac{1}{2}$ or spin $s_{z}=-\tfrac{1}{2}$. Let
+$\Pi_{\uparrow}=|\uparrow\rangle\langle\uparrow|=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right)\;,\qquad\Pi_{\downarrow}=|\downarrow\rangle\langle\downarrow|=\left(\begin{array}[]{cc}0&0\\ 0&1\end{array}\right)$ (3.5)
+be the projection operators on the spin up (down) states; the relative intensities of the upper and lower blots on the screen are, according to QM,
+${\bar{\Pi}_{\uparrow}}=\mathop{Tr}\nolimits\Pi_{\uparrow}\rho=\rho_{11}\;;\quad{\bar{\Pi}_{\downarrow}}=\mathop{Tr}\nolimits\Pi_{\downarrow}\rho=\rho_{22}\;,\qquad\rho_{11}\ +\rho_{22}=1\;.$ (3.6)
+The prediction about the relative intensities of the two narrow atomic image bands from the wave function (3.2) and from the density matrix, (3.6), is in general indistinguishable, as is well known.
+3.2 Large spin
+Consider now the state of a spin $j$ directed towards a direction ${\mathbf{n}}$, $|j,{\mathbf{n}}\rangle$, given in (2.15),(2.26), and (2.27), Fig.
+4.
+Choosing the magnetic field (and its gradients) in the ${z}$-direction, we need to express $|j,{\mathbf{n}}\rangle$ as a superposition of the eigenstates of ${\hat{J_{z}}}$ ($j=n/2$).
+This is precisely the expression given in Eq.(2.26) and Eq.(2.27).
+Using Stirling’s formula in (2.27), one finds, for $n$ and $k=j+m$ both large with $x=k/n$ fixed, the following distribution in different values of $m=-j+k$,
+$|c_{k}|^{2}\simeq e^{nf(x)}\;,\qquad x=k/n=(j+m)/2j\;,$ (3.7)
 where
-${\hat{J}}^{2}|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\;.$ (3.12)
-Actually, all states “almost oriented towards ${\mathbf{n}}$”, $|j,{\mathbf{n}},q\rangle,$
-$\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,,\,\,\,{\rm finite},$ (3.13)
-discussed in Sec.
-2.2.3 (see (2.22)), can also be interpreted as ${\bf j}^{(\rm class)}=j\,{\bf n}$.
-The aim of our present investigation is to examine the consistency of such a picture, with respect to the SG processes (this has already been done in Sec.
-2.2), the addition rule of the angular momenta (Sec.
-4 below) and the rotation matrix (Sec.
-5 below).
-As for the states with strong quantum fluctuations $|j,m\rangle$, (3.9) (with $m$ generic and not close to ±$\pm j$), they are certainly quantum mechanical, independently of $j$.
-The same can be said of a generic spin $j$ wave function, (3.4).
-If such a particle is sent into a SG set-up, it could split in an arbitrary distribution of sub wavepackets, in contrast to the cases discussed in Sec.
-2.2.
-This results in a macroscopic quantum system.
-It is possible that such a strongly fluctuating, macroscopic quantum angular-momentum state can be realized experimentally, if the whole system is brought to extremely low temperatures near $T=0$  [^6] We recall that any system is in its quantum-mechanical ground state at $T=0$.
-Accordingly many macroscopic systems can be brought to a quantum-mechanical state, if sufficiently low temperatures can be attained.
-See the references cited in [1] for the recent efforts to fabricate such systems experimentally..
-To see what happens at finite temperatures (either the body-temperature of the macroscopic body, or a warm environment) instead, it is useful to recall the so-called Born-Einstein dispute and its resolution (see [1]).
-Einstein strongly rebuked the idea by Born, that the absence of quantum diffusion is sufficient to explain the classical nature (the unique trajectoriy) of a macroscopic body, arguing that a doubly or multiply split wave packet, with their centers separated by macroscopic distances, are allowed by Schrödinger equations even for a macroscopic particle.
-The latter is certainly non-classical.
-The missing piece for solving this puzzle turns out to be the temperature [1].
-Even though at exceptionally low temperatures such a state is possible, it is not so at finite temperatures.
-Emission of photons and the ensuing self-decoherence (in the case of an isolated body) [1] or an environment-induced decoherence (for an open system) [5, 6, 7] makes a split system a mixture (a mixed state) essentially instantaneously.
-Also, they cannot be prepared experimentally, e.g., by a passage through a double slit [1].
-A (macroscopic) particle passes either through one or the other slit, due to absence of diffusion.
-A similar consideration applies to the general angular momentum states with strong fluctuations, (3.9) or (3.4).
-In the large $j$ limit, such a state necessarily involves large space extension: a macroscopic quantum state.
-A macroscopic pure quantum state can be realized only at exceedingly low temperatures  [^7] To tie up possible loose ends of the argument, let us note that a single, spin $1/2$ isolated atom in its ground state (a microscopic system) travelling in a good vacuum, has effectively $T=0$.
-Thus the fact that its split wave packets can attain a macroscopic extension - i.e.
-its being in a macroscopic quantum state - is perfectly consistent.
-To realize a macroscopic system (made of many atoms and molecules) in a pure quantum state is another story: it is much more difficult to prepare the necessary low temperatures close to $T=0$.
-.
-See [1] for more discussions.
-At finite temperatures, macroscopically split wave packet of such a particle under the SG set-up is necessarily an incoherent mixture.
-What is rather remarkable, perhaps, is the fact that the angular momentum states of minimum fluctuations, (3.12), (3.13), though pure, effectively become classical at $j\to\infty$, without need of any (thermal or environment-induced) decoherence effects, hence independently of the temperatures.
-This is in agreement with the known general semi-classical behavior of a (pure-state) wave function in QM in the limit $\hbar\to 0$.
-3.1 Orbital angular momentum
+$f(x)=-x\log x-(1-x)\log(1-x)+2x\log\cos\tfrac{\theta}{2}+2(1-x)\log\sin\tfrac{\theta}{2}\;.$ (3.8)
+The saddle-point approximation valid at $n\to\infty$, yields
+$f(x)\simeq-\frac{(x-x_{0})^{2}}{x_{0}(1-x_{0})}\;,\qquad x_{0}=\cos^{2}\tfrac{\theta}{2}\;,$ (3.9)
+and therefore
+[…]
+in the $n\to\infty$ ($x=k/n$ fixed) limit. The narrow peak position $x=x_{0}$ corresponds to (see Eq.(2.26))
+$J_{z}=m=n(x-\tfrac{1}{2})=j\,(2\cos^{2}\tfrac{\theta}{2}-1)=j\,\cos\theta\;.$ (3.11)
+This means that a large spin ($j\gg\hbar$) quantum particle with spin directed towards ${\mathbf{n}}$, in a Stern-Gerlach setting with an inhomogeneous magnetic field (3.3), moves along a single trajectory of a classical particle with $J_{z}=j\,\cos\theta$, instead of spreading over a wide range of split sub-packet trajectories covering $-j\leq m\leq j$.
+See Fig.
+5 (spin $j=10^{3}$) and Fig.
+6 ($j=2\cdot 10^{5}$) and compare them with Fig.
+1 for spin $j=13/2$.
+Figure 5: The distribution $|c_{k}|^{2}$ in $k$ for possible values of $J_{z}=m$ ($m=-j+k$), for a spin $j=1\cdot 10^{3}$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange peak) or with $\theta=\pi/4$ (right, blue peak).
+The particle starts looking like classical.
+Figure 6: The distribution $|c_{k}|^{2}$ ($J_{z}=m=-j+k$) as in Fig.
+5 but for a spin $j=2\cdot 10^{5}$, for $\theta=\pi/2$ (center, orange peak) or $\theta=\pi/4$ (right, blue peak).
+This figure is drawn by using the approximation (3.7) and (3.8), rather than the exact formula, (2.26), (2.27).
+This (perhaps) somewhat surprising result appears to indicate that quantum mechanics (QM) takes care of itself, so to speak, in ensuring that a large spin particle ($j/\hbar\to\infty$) behaves classically, at least for these particular states $|j,{\mathbf{n}}\rangle$.
+No extra conditions are necessary.
+This is consistent with the known general behavior of the wave function in the semi-classical limit ($\hbar\to 0$)  [^5] Of course this does not mean that the classical limit necessarily requires or implies $j/\hbar\to\infty$..
+Note that if the value $j\gg\hbar$ is understood as due to the large number of spin $\frac{1}{2}$ particles composing it (see the comment after (2.27)), the spike (3.9), (3.10) can be understood as due to the accumulation of a huge number of microstates giving $J_{z}\simeq m$.
+Another, equivalent way to see the shrinkage of the distribution in $J_{z}=m$, (3.10), is to study its dispersion.
+Given the relative weights $|c_{k}|^{2}$, (2.27), where $m=-j+k$ and $j=\tfrac{n}{2}$, with $k=0,1,\ldots,n$, it is a simple exercise to find that
+$\langle J_{z}\rangle=j\,\cos\theta\;,\qquad\langle J_{z}^{2}\rangle=j^{2}\cos^{2}\theta+\frac{j}{2}\sin^{2}\theta\;,$ (3.12)
+so that the dispersion (the fluctuation width) is given by  [^6] These results are well known [10].
+$(\Delta J_{z})^{2}=\langle(J_{z}-\langle J_{z}\rangle)^{2}\rangle=\langle J_{z}^{2}\rangle-\langle J_{z}\rangle^{2}=\frac{j}{2}\sin^{2}\theta\;.$ (3.13)
+It follows that
+$\lim_{j/\hbar\to\infty}\frac{|\Delta J_{z}|}{j}=0\;,$ (3.14)
+i.e.
+it is an infinitely narrow distribution.
+Note that the SG experiment is indeed a measurement of $J_{z}$ in the state (2.26).
+The fact that the result is always $J_{z}\sim\langle J_{z}\rangle=j\,\cos\theta$ with no dispersion, means that it is a classical angular momentum $\sim j\,{\mathbf{n}}$.
+3.2.1 The state preparation and successive SG set-ups
+For a small $j$, as it is well known, the SG set-up can be used to prepare any desired state $|j,J_{z}=m\rangle$, by an intentionally biased SG apparatus, e.g.
+in which all the split sub-wavepackets are blocked, except for the one corresponding to $|j,m\rangle$.
+Equivalently, one may use the so-called null-measurements (see Ref.
+[30] for a recent review).
+In any case, for a small finite $j$, where there are $12j+$ 1 well-separated sub-packets (see Fig.
+1), there is no difficulty in extracting any desired quantum state for subsequent experiments (state preparation).
+In the large-spin limit ($j\to\infty$) of the Bloch states $|j,{\mathbf{n}}\rangle$ we are considering, there is no possibility of extracting or selecting a desired sub-wavepacket, since there is essentially only one narrow wavepacket present (see (3.11), Fig.
+6).
+Trying to extract a state with $J_{z}$ not close to $j\cos\theta$ would simply give a zero result.
+We have in mind a macroscopic scenario where, in comparison to the finite region (in principle) occupied by the waves $-j\leq m\leq j$, the fluctuations of the values of $m$ around (3.11) are of the order of $\sqrt{j}\ll j$, i.e.
+of zero width.
+We might consider introducing a second set of Stern-Gerlach (SG) magnetic fields directed in a different direction, denoted by ${\hat{z}}^{\prime}$, as a new quantization axis.
+This setup allows us to study the state emerging from the first set of SG magnetic fields (but without an imaging screen),
+$|j,{\mathbf{n}}\rangle\sim c_{k}\,|j,J_{z}=m\rangle\;,\qquad m\sim j\,\cos\theta\;.$ (3.15)
+Since the sum of states around $m\sim j\,\cos\theta$ is essentially equal to the complete sum, this state can be rewritten as
+$|j,{\mathbf{n}}\rangle$ $=$ $\sum_{m,m^{\prime}}c_{k}\,|j,m^{\prime}\rangle\langle j,m^{\prime}|j,m\rangle$ (3.16)
+$=$ $\sum_{m,m^{\prime}}|j,m^{\prime}\rangle\langle j,m^{\prime}|j,m\rangle\langle j,m|j,{\mathbf{n}}\rangle$
+$=$ $\sum_{m^{\prime}}|j,m^{\prime}\rangle\langle j,m^{\prime}|j,{\mathbf{n}}\rangle=\sum_{m^{\prime}}c_{k^{\prime}}|j,m^{\prime}\rangle\;,$
+where in the first line a completeness relation $\sum_{m^{\prime}}|j,m^{\prime}\rangle\langle j,m^{\prime}|={\mathbf{1}}$ has been inserted, and in the second we recalled $c_{k}=\langle j,m|j,{\mathbf{n}}\rangle$ (see Eq.(2.26)).
+(3.16) represents an expansion of the original state $|j,{\mathbf{n}}\rangle$ as a linear combination of the eigenstates of $J_{z^{\prime}}=m^{\prime}$.
+In the limit $j\to\infty$, the results from Section 3.2 tell us that the sum is dominated by the small region around $m^{\prime}\sim j\cos({\mathbf{n}}\cdot{\hat{z}}^{\prime})$, with negligible fluctuations.
+The trajectory remains unique, regardless of the choice of quantization axis (i.e.
+the direction of the magnetic field) ${\hat{z}}^{\prime}$.
+3.2.2 States almost directed towards $\vec{\mathbf{n}}$
+Obviously, in the large spin limit, other states close to $|j,{\mathbf{n}}\rangle$,
+$\left({\hat{\mathbf{j}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,\,\,{\rm fixed}$ (3.17)
+(i.e. the spin roughly oriented towards ${\mathbf{n}}$) will behave similarly. For instance, the $q=1$ state is (cfr. Eq.(2.26))
+$|j,{\mathbf{n}},1\rangle$ $=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}\,|{\bf n}\rangle_{1}\otimes\cdots|{\bf n}\rangle_{\ell}\otimes|-{\bf n}\rangle\otimes|{\bf n}\rangle_{\ell+2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (3.18)
+$=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}|j_{1},{\mathbf{n}}\rangle\otimes|-{\bf n}\rangle\otimes|j_{2},{\mathbf{n}}\rangle\;,$
+where $|j_{1},{\mathbf{n}}\rangle$ and $|j_{2},{\mathbf{n}}\rangle$ ($j_{1}=\tfrac{\ell}{2}\;,j_{2}=\tfrac{N-\ell-1}{2}$), are two spins oriented towards ${\mathbf{n}}$, (2.15), as already studied in Sec.
+3.2.
+In the $j\to\infty$ limit, the sum is dominated by terms with $\ell\sim\mathcal{O}(N)\to\infty$.
+The SG projection of this state onto the $j_{z}$ eigenstates is therefore approximated by
+$j_{z}\sim j_{1\,z}+j_{2\,z}=j_{1}\cos\theta+j_{2}\cos\theta\simeq j\,\cos\theta\;,$ (3.19)
+where the results of Sec.
+3.2 have been used.
+A similar result holds for $|j,{\mathbf{n}},q\rangle$ for any finite $q$, showing that all these states behave as a classical angular momentum, $j\,{\mathbf{n}}$.
+4 Orbital angular momentum
 All our formal arguments (2.11)–(2.17) apply equally well, both to spin (the intrinsic angular momentum of a particle or the total angular momentum of the center of mass of a body in its rest frame) and to orbital angular momentum.
 However, our primary tool of analysis is the Stern-Gerlach (SG) set-up (Sec.
-2.2), which is more suitable for spin than for an angular momentum associated with the orbital motion of a particle.
+3.2), which is more suitable for spin than for an angular momentum associated with the orbital motion of a particle.
 One might wonder how the large angular momentum limit works for the latter.
@@
 where $P_{\ell}^{m}(x)$ are the associated Legendre polynomials.
-Let us consider the state of minimum fluctuation, $|\ell,\ell\rangle$ (see Eq.
-(3.10)).
+Let us consider the state of minimum fluctuation, $|\ell,\ell\rangle$ (see Eq.(2.14)).
 It is easy to see that the angular distribution is strongly peaked at $\theta=\tfrac{\pi}{2}$ at large $\ell$,
@@
 8 for $\ell=m=10^{3}$.
-In states $|\ell,m\rangle$ with stronger fluctuations ($m$ not close to $\ell$), the distribution instead covers a large portion of the $3$D solid angle, for any $\ell$ (see Fig.
+In states $|\ell,m\rangle$ with stronger fluctuations, with $m$ not close to $\pm\ell$, the distribution instead covers a large portion of the $3$D solid angle, for any $\ell$ (see Fig.
 8 for $\ell=100,m=50$).
@@
 5 Addition of angular momenta
-Let us now consider two spins, ${\mathbf{j}}_{1}$ and ${\mathbf{j}}_{2}$. The composition-decomposition rules in QM is well known, i.e.
+Let us now consider two spins, ${\mathbf{j}}_{1}$ and ${\mathbf{j}}_{2}$. The composition-decomposition rule in QM is well known, i.e.
 ${\underline{j_{1}}}\otimes{\underline{j_{2}}}={\underline{j_{1}+j_{2}}}\oplus{\underline{j_{1}+j_{2}-1}}\oplus\ldots{\underline{{|j_{1}-j_{2}|}}}\;,$ (5.1)
-where $\underline{j_{1}}$, $\underline{j_{2}}$, indicates the multiplet (the irreducible representation) of SU(2)  [^8] We recall that SU(2) is the simply-connected double cover of the orthogonal group SO(3)., with multiplicity $12j+$ 1 and with $\max\{j_{z}\}=j$.
+where $\underline{j}$ indicate the multiplet (the irreducible representation) of SU(2)  [^7] We recall that SU(2) is the simply-connected double cover of the orthogonal group SO(3)., with multiplicity $12j+$ 1 and with $\max\{j_{z}\}=j$.
 We set $\hbar=1$ here.
 We wish to find out how the addition rule looks like in the limit $j_{1},j_{2}\to\infty$.
-For the reasons discussed in Sec.
-3, we are particularly interested in the composition rule for angular momentum states with minimum fluctuations, of the form, $|j,{\mathbf{n}}\rangle$.
+We are particularly interested in the composition rule for angular momentum states with minimum fluctuations, of the form, $|j,{\mathbf{n}}\rangle$.
 Therefore, let us consider two particles (spins) in the states,
@@
 $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\;,$ (5.3)
-in the large $j_{1},j_{2}$ limit. Since the choice of the axes is arbitrary, we may take
+in the large $j_{1},j_{2}$ limit.
+Since the choice of the axes is arbitrary, we may take
 ${\mathbf{n}}_{2}=(0,0,1)\;.$ (5.4)
@@
 where $|j_{2},m_{2}=j_{2}\rangle$ is the highest $\hat{j}_{2\,z}$ state.
-We added the tilde sign on ${\mathbf{n}}_{1}$ to indicate that it is the vector ${\mathbf{n}}_{1}$ in the reference system of (4.4).
+We added the tilde sign on ${\mathbf{n}}_{1}$ to indicate that it is the vector ${\mathbf{n}}_{1}$ in the reference system, (5.4).
 Now $|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle$ in the $j_{1}\to\infty$ limit has been studied in Sec.
-2.
+3.2.
 It is
-$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle=\sum_{m=-j_{1}}^{j_{1}}c_{m}\,|j_{1\,z}=m\rangle\;,\qquad$ (4.6)
+$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle=\sum_{m=-j_{1}}^{j_{1}}c_{m}\,|j_{1},j_{1\,z}=m\rangle\;,$ (5.6)
 where the sum is dominated by the values of $m$ around
-$m=j_{1\,z}\simeq j_{1}\cos\theta\;,\qquad\cos\theta={\tilde{\mathbf{n}}}_{1\,z}={\mathbf{n}}_{1}\cdot{\mathbf{n}}_{2}\;.$ (4.7)
-Since the eigenvalues of $\hat{j_{z}}$ simply add up in the product state, the expansion of $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle$ in the eigenstates of $\hat{j}_{{\rm tot}\,z}$ is dominated by terms
-$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{{\tilde{n}}_{2}}\simeq j_{1}\cos\theta+j_{2}\rangle\;.$ (4.8)
+$m=j_{1\,z}=j_{1}\cos\theta\;,\qquad\cos\theta={\tilde{\mathbf{n}}}_{1\,z}={\mathbf{n}}_{1}\cdot{\mathbf{n}}_{2}\;,$ (5.7)
+with the fluctuations of $j_{1\,z}$ which become negligible in the infinite spin limit (see Fig. 6). For the purpose of the discussion of this section, we introduce the notation
+$|j_{1},{\tilde{\mathbf{n}}}_{1}\rangle\simeq|j_{1},j_{1\,z}=j_{1}\cos\theta\rangle\;,$ (5.8)
+by using the $\simeq$ sign, to express this fact.
+Since the eigenvalues of $\hat{j_{z}}$ simply add up in the product state, the expansion of $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle$ in the expansion in terms of the eigenstates of $\hat{j}_{{\rm tot}\,z}$ is dominated by terms with $j_{\rm tot}^{{\tilde{n}}_{2}}=j_{1}\cos\theta+j_{2}$, namely,
+$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{{\tilde{n}}_{2}}=j_{1}\cos\theta+j_{2}\rangle\;,$ (5.9)
+where $j_{\rm tot}^{{\tilde{n}}_{2}}$ means the component in the direction of ${\tilde{n}}_{2}$ of the (for the moment, unknown) total spin $j_{\rm tot}$.
 Exchanging ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$ and repeating the arguments, one finds also that
-$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{\tilde{n}_{1}}\simeq j_{1}+j_{2}\cos\theta\rangle\;.$ (4.9)
-But which quantum angular-momentum state does the product state (4.3) represent? The answer is that
+$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},j_{\rm tot}^{\tilde{n}_{1}}=j_{1}+j_{2}\cos\theta\rangle\;.$ (5.10)
+But which quantum angular-momentum state does the product state (5.3) represent? Such a state must be compatible with the projections (5.9) and (5.10). The answer is that
 $|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot},{\mathbf{n}}\rangle\;,$ (5.11)
-where $j_{\rm tot}$ and the unit vector ${\mathbf{n}}$ are such that (i.e. defined by)
+with $j_{\rm tot}$ and the unit vector ${\mathbf{n}}$ defined by
 $j_{\rm tot}\,{\mathbf{n}}=j_{1}{\mathbf{n}}_{1}+j_{2}{\mathbf{n}}_{2}\;.$ (5.12)
-The proof of this is as follows.
+Note that this (classical) vector sum determines both $j_{\rm tot}$ (the magnitude) and ${\mathbf{n}}$ (the direction) uniquely.
+The quantum state $|j_{\rm tot},{\mathbf{n}}\rangle$ is defined via the correspondence (2.15), (2.16) proposed before.
+The proof of the statement (5.11) is as follows.
 By selecting the directions of the SG magnets directions (the quantization axis) to ${\mathbf{n}}_{2}$ or to ${\mathbf{n}}_{1}$ on the state $|j_{\rm tot},{\mathbf{n}}\rangle$, (5.12), one can apply the results of Sec.
 3.1 to get exactly the results (5.9) or (5.10), respectively.
-Now, according to the quantum-classical correspondence (Eq.
-(3.11) and Eq.
-(3.12)), the state $|j_{\rm tot},{\mathbf{n}}\rangle$ (see Eq.
-(4.11)) translates to
+Now, according to the quantum-classical correspondence of Eq.(2.16), the state $|j_{\rm tot},{\mathbf{n}}\rangle$ (see Eq.(5.12)) translates into
 ${\mathbf{j}}_{\rm tot}^{(\rm class)}=j_{1}\,{\mathbf{n}}_{1}+j_{2}\,{\mathbf{n}}_{2}\;.$ (5.13)
-But, in view of (3.11), this is precisely the addition rule of the two classical angular momenta (i.e. a vector sum).
-To be complete, one must check that the projection on any generic direction ${\mathbf{n}}_{3}$ works correctly.
+But, in view of (2.16), this is precisely the addition rule of the two classical angular momenta.
+To be complete, one must check that the projection on any other generic direction ${\mathbf{n}}_{3}$ works correctly.
 By using the result of Sec.
-2 for the SG projection on a new ${\mathbf{n}}_{3}$ axis, one finds
-$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot}^{\prime},j_{\rm tot}^{\prime\,n_{3}}\simeq j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}\rangle\;,$ (4.13)
+3.2 for the SG projection on a new ${\mathbf{n}}_{3}$ axis, one finds
+$|j_{1},{\mathbf{n}}_{1}\rangle\otimes|j_{2},{\mathbf{n}}_{2}\rangle\simeq|j_{\rm tot}^{\prime},j_{\rm tot}^{\prime\,n_{3}}=j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}\rangle\;,$ (5.14)
 where we made use of the fact that ${\hat{\mathbf{j}}}_{1}$ and ${\hat{\mathbf{j}}}_{2}$ commute and the eigenvalues of their ${\mathbf{n}}_{3}$ components just sum up in the direct product state.
 Also, to be conservative, we have left $j_{\rm tot}^{\prime}$ unknown.
-However,
+But
 $j_{1}{\mathbf{n}}_{1}\cdot{\mathbf{n}}_{3}+j_{2}{\mathbf{n}}_{2}\cdot{\mathbf{n}}_{3}=(j_{1}{\mathbf{n}}_{1}+j_{2}{\mathbf{n}}_{2})\cdot{\mathbf{n}}_{3}\;,$ (5.15)
-thus
+therefore
 ${\mathbf{j}}_{\rm tot}^{\prime}={\mathbf{j}}_{\rm tot}=j_{1}\,{\mathbf{n}}_{1}+j_{2}\,{\mathbf{n}}_{2}\;,$ (5.16)
 by again using the results of Sec.
-2.1 in Eq.
-(4.13).
+3.1 in Eq.(5.14), which proves Eq.(5.11) and Eq.(5.12).
 In conclusion, the direct product state, (5.2), (5.3), has a simple classical interpretation.
-It corresponds to the classical sum (the vector addition) of two angular momenta.
+It corresponds to the classical sum (the vector addition) of two angular momenta, see Fig.
+9.
+Figure 9: The direct product state, (5.3), behaves as a classical vector sum, in the limit, $j_{1},j_{2}\to\infty$, in the sense of the correspondence, (2.16).
 From the perspective of the general composition rule of two angular momentum states (5.1), what we have seen are the results concerning a specific pair of states (i.e.
 those of minimum fluctuations) in ${\underline{j_{1}}}$ and ${\underline{j_{2}}}$, characterized by the “orientations” ${\bf n}_{1}$ and ${\bf n}_{2}$.
-The total angular momentum magnitude in the product state (4.11) satisfies
+The total angular momentum magnitude in the product state (5.11), (5.12) satisfies
 $j_{1}+j_{2}\geq j_{\rm tot}=|j_{1}{\mathbf{n}}_{1}+j_{2}{\mathbf{n}}_{2}|\geq|j_{1}-j_{2}|\;,$ (5.17)
-depending on the relative orientation of ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$. This is consistent with the quantum-mechanical addition law, (4.1).
+depending on the relative orientation of ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$. Eq.(5.17) is nicely consistent with the quantum-mechanical composition law (5.1).
 6 Rotation matrix
-The rotation matrix for a general spin $j$ can be constructed by taking the direct products of the rotation matrix for spin $1/2$, i.e.
+The rotation matrix for a general spin $j$ can be constructed by taking the direct products of the rotation matrix for spin $\frac{1}{2}$, i.e.
 (2.7).
 This is based on the fact that any angular momentum state, $|j,m\rangle$, can be constructed as the system made of $N$ spin $\tfrac{1}{2}$ particles, completely symmetric under exchange of the particles, a fact already conveniently used in Sec.
-2.1.
-The theory of angular momentum can be fully reconstructed this way [8] by using the operators,
-${\hat{J}}^{a}=\sum_{i,j=1,2}a_{i}^{\dagger}\Big{(}\frac{\tau^{a}}{2}\Big{)}_{ij}a_{j}\;,\qquad[{\hat{J}}^{a},{\hat{J}}^{b}]=i\,\epsilon_{abc}{\hat{J}}^{c}\;,$ (5.1)
-where $\tau^{a}$ are the Pauli matrices, $a_{i},a^{\dagger}_{i}$ are the harmonic-oscillator annihilation and creation operators, of spin-up ($i=1$) and spin-down ($i=2$) particles,
-$[a_{i},a^{\dagger}_{j}]=\delta_{ij}\;.$ (5.2)
-[…]
-The rotation matrix for a generic spin $j$ is simply (writing the Euler angles symbolically as $\Omega\equiv\{\gamma,\theta,\phi\}$) the direct productof the rotation matrices for spin $1/2$, (3.3),
+3.1.
+The rotation matrix for a generic spin $j$ is simply (writing the Euler angles symbolically as $\Omega\equiv\{\gamma,\theta,\phi\}$) the direct productof the rotation matrices for spin $\frac{1}{2}$, (2.7),
 $(D^{j}(\Omega))_{M,M^{\prime}}={\cal S}\left[(D^{1/2}(\Omega))_{m_{1},m_{1}^{\prime}}(D^{1/2}(\Omega))_{m_{2},m_{2}^{\prime}}\cdots(D^{1/2}(\Omega))_{m_{N},m_{N}^{\prime}}\right]$ (6.1)
@@
 $|j,{\mathbf{n}}\rangle=|\tfrac{1}{2},{\mathbf{n}}\rangle\otimes|\tfrac{1}{2},{\mathbf{n}}\rangle\otimes\ldots|\tfrac{1}{2},{\mathbf{n}}\rangle\;,$ (6.4)
-where $|\tfrac{1}{2},{\mathbf{n}}\rangle$ is given in Eq.
-(2.2), i.e.
-on the states with all spin $\tfrac{1}{2}$’s oriented in the same direction, is very simple.
-It is just the operation of rotating each and all spins towards the ${\hat{z}}$ direction (setting $\gamma=0$),
-$(D^{j}(\Omega))|j,{\mathbf{n}}\rangle=|\!\uparrow\rangle|\!\uparrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle=(1,0,\ldots,0)^{T}=|j,{\hat{z}}\rangle\;.$ (5.7)
+where $|\tfrac{1}{2},{\mathbf{n}}\rangle$ is given in Eq.(3.2), i.e.
+on the states with all spin $\tfrac{1}{2}$’s oriented in the same direction, is particularly simple.
+It is just the operation of rotating each and all spins towards the ${\hat{z}}$ direction (setting $\gamma=0$), see (2.6),
+$(D^{j}(\Omega))|j,{\mathbf{n}}\rangle=|\!\uparrow\rangle|\!\uparrow\rangle|\!\uparrow\rangle\cdots|\!\uparrow\rangle=(1,0,\ldots,0)^{T}=|j,j\rangle\;.$ (6.5)
 This corresponds precisely to the rotation of a vector,
@@
 where $R(\Omega)$ is the standard $3\times 3$ rotation matrix of classical mechanics.
+The discussion of this section is just a simple way of reproducing the well-known fact that the coherent spin (Bloch) states $|j,{\mathbf{n}}\rangle$ can be obtained by applying the rotation operator on the state pointing towards ${\hat{z}}$, $|j,j_{z}=j\rangle$ [5, 6, 7, 10, 11].
+7 General large-spin states
+In Sec.
+3 $\sim$ Sec.
+6 we have seen how the spin coherent (Bloch) states, (2.16), behave effectively as a classical angular momentum, $j{\mathbf{n}}$, in the $j\to\infty$ limit.
+It is an entirely different story with a generic large spin $j$ state in ${\mathbf{CP}}^{2j}$,
+$|j\,\,...\rangle=\sum_{m=-j}^{j}b_{m}\,|j,m\rangle\;,\qquad\sum_{m=-j}^{j}|b_{m}|^{2}=1\;.$ (7.1)
+When a particle carrying such a spin enters a Stern-Gerlach set-up with a strong inhomogeneous magnetic field, its wave function will in general split in many subwavepackets.
+See Fig.
+10 for several arbitrarily chosen distributions $\{b_{m}\}$, to be compared with Fig.
+5 or Fig.
+6.
+Figure 10: $|b_{m}|^{2}$, $J_{z}=m$, for a spin $j=1\cdot 10^{3}$ particle in the state (7.1), with four randomly chosen sets of coefficients $\{b_{m}\}$.
+To discuss the physics of such states, it is useful to recall the so-called Born-Einstein dispute (see e.g., [31]).
+Einstein strongly rebuked the idea by Born that the absence of quantum diffusion should be sufficient to explain the classical nature (the unique trajectoriy) of a macroscopic body, by saying that a doubly or multiply split wave packet, with their centers separated by a macroscopic distance, are allowed by Schrödinger equations, even for a macroscopic particle.
+Such a split wave function is certainly non-classical.
+The missing piece for solving this apparent puzzle turns out to be the temperature [22].
+Even though at exceptionally low temperatures such a state is certainly possible, it is not so at finite temperatures.
+Emission of photons and the ensuing self-decoherence (in the case of an isolated body with finite body temperatures) [22] or an environment-induced decoherence (for an open system) [31, 32, 33], makes a split system a mixture (a mixed state) essentially instantaneously.
+Also, they cannot be prepared experimentally, e.g., by a passage through a double slit [22].
+A (macroscopic) particle passes either through one or the other slit, due to absence of diffusion and/or due to decoherence.
+The particle after the passage is behind either one or the other slit, even without any measurement.
+The status of general angular momentum states with strong fluctuations, (7.1), is similar to that of Einstein’s macroscopic split wave packet.
+In the large $j$ limit, a generic large spin state (represented by randomly distributed stars, Fig.
+3), far from the Bloch state $|j,\mathbf{n}\rangle$, necessarily acquires a large space extension under an inhomogeneous magnetic field: a macroscopic quantum state.
+See Fig.
+10.
+But a macroscopic pure quantum state is possible only at exceedingly low temperature close to $T=0$: otherwise, it is a mixture.
+See [22] for more discussions.
+To tie up possible loose ends of the discussion, note that a single, spin $\frac{1}{2}$ isolated atom in its ground state (a microscopic system) travelling in a good vacuum, is effectively a system at $T=0$.
+Thus the fact that its split wave packets can get separated by a macroscopic distance in an SG process - i.e.
+its being in a macroscopic quantum state - is perfectly consistent with the general argument above.
+To realize a macroscopic system composed of many atoms and molecules in a pure quantum state, is another story: it is much more difficult to prepare the necessary low temperatures close to $T=0$, to maintain its pure-state nature.
+At finite temperatures, a macroscopically split wave packet of such a particle, which might arise under the SG set-up, as in Fig.
+10, is necessarily an incoherent mixture.
+An important point to keep in mind, however, is the following.
+The fact that a macroscopic quantum state such as the split wavepackets of Fig.
+10 becomes a mixture under the decoherence effects at finite temperatures [31, 32, 33], does not mean that it becomes classical.
+Decoherence and classical limits are two distinct phenomena in general [23].
+This point can hardly be overemphasized.
+What is indeed remarkable, perhaps, is the fact that the angular momentum states of minimum fluctuations, (1.4), (2.15), (2.26), (2.27), though pure, effectively become classical at $j\to\infty$, without help of any (thermal or environment-induced) decoherence effects, hence independently of the temperatures.
+This might appear to be in line with the familiar discussion on the semi-classical limit of a (pure-state) wave function in QM at $\hbar\to 0$, but, our discussion about the generic large spin states indicates the presence of a loophole in the argument that only the dimensionless ratio $j/\hbar$ should matter.
 8 Conclusion
-The observations made in this note render the concept that a spin ($S$) (or angular momentum, $j$) becomes classical in the limit, $S(j)\to\infty$, a more precise one.
-In particular, we found the identification between the states with minimum fluctuations ((3.11), (3.12), and (3.13)) with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of SG experiments, of the angular-momentum addition rule, and of the rotation matrix.
-At the same time, our analysis has revealed a subtlety in the quantum-classical correspondence in this context.
+The observations made in this note render the concept that a spin (or angular momentum), $j$, becomes classical in the limit, $j\to\infty$, a more precise one.
+In particular, we found the identification between the states with minimum fluctuations (2.15) - the coherent spin states - with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of the SG experiments, with the angular-momentum composition rule, and with the rotation matrix.
+At the same time, our analysis has revealed a subtlety in the quantum-classical correspondence in the large spin limit in general.
 It reflects the difference between the spaces of the quantum-mechanical and classical angular momentum states of definite magnitude ($\mathbf{CP}^{2j}$ and $\mathbf{CP}^{1}\sim S^{2}$, respectively).
-It is the angular momentum states of minimum fluctuations, (3.11), (3.12), (3.13), that naturally smoothly transit into classical angular momentum in the $S(j)\to\infty$ limit, independently of temperatures.
-On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ are possible as pure quantum states only at exceedingly low temperatures.
+It is the angular momentum states of minimum fluctuations, (1.4),(2.15),(2.26),(2.27), that naturally smoothly become classical angular momentum in the $j\to\infty$ limit, independently of the temperatures.
+On the other hand, more general, strongly fluctuating angular momentum states belonging to $\mathbf{CP}^{2j}$ remain quantum-mechanical even in the $j\to\infty$ limit.
+They can stay pure quantum states only at exceedingly low temperatures, and isolated.
 The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other.
-A macroscopic system is necessarily in a mixed state at finite temperatures [5, 6, 7],[1].
-For instance, no macroscopically split pure wave packets in a SG setting survive decoherence effects.
-The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles seems to be a fruitful one both mathematically and physically.
-From the mathematics point of view, the entire theory of angular momentum can be reconstructed this way [8], as we recalled already, and it helps to recover certain formulas easily.
-It looks also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized model of a macroscopic body, made of many atoms and molecules (spins).
-The sharp spike of the SG projection of the state $|S,{\bf n}\rangle$ in the narrow region around $S_{z}=S\cos\theta$ can then be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $S_{z}$.
-The various form of suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is in fact analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [9], to explain the exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones.
+A macroscopic system is in a mixed state at finite temperatures [31, 32, 33],[22].
+No macroscopically split pure wave packets in a SG setting survive decoherence.
+At the same time, as pointed out in [23], a macroscopic quantum state which becomes a mixture due to decoherence, remains such: a mixed quantum state.
+Decoherence in itself does not make the system considered classical, contrary to the statements often made in the literature.
+The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles turned out to be quite a fruitful one both mathematically and physically, as seen in several occasions.
+From the mathematical point of view, the entire theory of angular momentum can be reconstructed this way [26], as we recalled already, and it helps to recover certain formulas easily.
+It is also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized (toy) model of a macroscopic body, made of many atoms and molecules (spins).
+Seen this way, the sharp spike of the SG projection of the state $|j,{\bf n}\rangle$ in the narrow region around $J_{z}=j\cos\theta$, Fig.
+5, Fig.
+6, can be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $J_{z}$.
+[^8] The suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is somewhat analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [34], to explain the general exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones.
+This being so, let us not forget that the deepest insight of Schrödinger in his book was that many biological processes such as mutation, reproduction, and inheritance, involve quantum mechanics in essential ways.
+Nevertheless, logically our discussion is about the quantum and classical properties of a single large spin, and is not about the thermodynamical or statistical properties of many body systems, as in [6, 7], for instance.
+And in spite of inevitable partial overlaps with numerous observations made in the literature [1]-[21], the ideas around the proposal (2.16) exposed in Sec.
+2, Sec.
+3, Sec.
+5, and Sec.
+7 are, to the best of our knowledge, new.
 Acknowledgments
-The work by KK is supported by the INFN special-initiative project grant GAST (Gauge and String Theories).
-We are grateful to Hans-Thomas Elze, Jarah Evslin, Riccardo Guida, Pietro Menotti and Arkady Vainshtein for useful comments and precious advices.
+The work by K.K.
+is supported by the INFN special-initiative project grant GAST (Gauge and String Theories).
+We are grateful to Hans Thomas Elze, Jarah Evslin, Riccardo Guida, Pietro Menotti and Arkady Vainshtein for useful comments and precious advices.
 Appendix A Newton’s equation for a macroscopic body
-The conditions needed for the CM of an isolated macroscopic body at finite body temperatures to obey Newton’s equations have been discussed in great care by one of the present authors [1]. They are
+The conditions needed for the CM of an isolated macroscopic body at finite body temperatures to obey Newton’s equations have been investigated in great care by one of the present authors [22].
+They are
   • (i)
@@
 [22] for various subtleties and for the explicit derivation of Newton’s equation under external gravitational forces, under weak, static, smoothly varying external electromagnetic fields, and under a harmonic-oscillator potential.
-Somewhat unexpectedly, the environment-induced decoherence [5, 6, 7] which is extremely effective in rendering macroscopic states in a finite-temperature environment a mixture, has been found not to be the most essential elements for the derivation of classical mechanics from quantum mechanics [1, 2].
+Somewhat unexpectedly, the environment-induced decoherence [31, 32, 33] which is extremely effective in rendering macroscopic states in a finite-temperature environment a mixture, is found not to be the most essential element for the derivation of classical mechanics from quantum mechanics [22, 23].
+Appendix B A subtle face of the Stern-Gerlach experiment
+The Hamiltonian is given by
+$H=\frac{{\bf p}^{2}}{2m}+V\;,\qquad V=-{\boldsymbol{\mu}\cdot{\bf B}}\;,$ (B.1)
+${\boldsymbol{\mu}}=\mu_{B}\,g\,{\mathbf{s}}\;,\qquad\partial B_{z}/\partial z\neq 0\;,$ (B.2)
+where $\mu_{B}=\frac{e\hbar}{2m_{e}c}$ is the Bohr magneton.
+We recall the well-known fact that the gyromagnetic ratio $g\simeq 2$ of the electron and the spin magnitude $1/2$ approximately cancel, so $\mu_{B}$ is the magnetic moment  [^9] For illustration purpose, we use this value below, but the discussion is obviously valid for general atoms., in the case of the atoms such as $A_{g}$, where a single outmost electron provides the total spin $\tfrac{1}{2}$.
+An example of the inhomogeneous field ${\bf B}$ appropriate for the Stern-Gerlach experiment is [29, 28]
+${\bf B}=(0,B_{y},B_{z}),\qquad B_{y}=-b_{0}\,y,\quad B_{z}=B_{0}+b_{0}\,z\;$ (B.3)
+which satisfy $\nabla\cdot{\bf B}=\nabla\times{\bf B}=0$. The constant field $B_{0}$ in the $z$ direction must be large,
+$|B_{0}|\gg|b_{0}\,y|\;.$ (B.4)
+in the relevant region of $(y,z)$ of the experiment. The wave function of the spin $\frac{1}{2}$ particle entering the SG magnet has the form,
+$\Psi={\tilde{\psi}}_{1}({\bf r},t)|\!\uparrow\rangle+\,{\tilde{\psi}}_{2}({\bf r},t)|\!\downarrow\rangle\;.$ (B.5)
+obeying the Schrödinger equation,
+$i\hbar\frac{d}{dt}\Psi=H\,\Psi\;.$ (B.6)
+By redefining the wave functions for the upper and down spin components as
+${\tilde{\psi}}_{1}({\bf r},t)=e^{i\mu_{B}B_{0}t/\hbar}\psi_{1}({\bf r},t)\;,\qquad{\tilde{\psi}}_{2}({\bf r},t)=e^{-i\mu_{B}B_{0}t/\hbar}\psi_{2}({\bf r},t)\;,$ (B.7)
+one finds that the up- and down- spin components $\psi_{1}$ and $\psi_{2}$ satisfy the separate Schrödinger equations [29]
+$i\hbar\frac{\partial}{\partial t}\psi_{1}=\left(\frac{{\bf p}^{2}}{2m}-\mu_{B}b_{0}z\right)\psi_{1}\;,\qquad i\hbar\frac{\partial}{\partial t}\psi_{2}=\left(\frac{{\bf p}^{2}}{2m}+\mu_{B}b_{0}z\right)\psi_{2}\;.$ (B.8)
+This is because the term $\propto-\mu_{y}B_{y}=-(g\mu_{B}b)s_{y}$ in the Hamiltonian (B.1) mixing the two components $\psi_{1,2}$ has acquired a rapidly oscillating phase factor,
+±$\pm i\mu_{B}\,b_{0}y\,e^{\mp 2i\mu_{B}B_{0}t/\hbar}\;,$ (B.9)
+hence can be safely neglected. The condition (B.4) is crucial here.
+As explained in [28], this can be classically understood as the spin precession effect around the large constant magnetic field $B_{0}\hat{z}$, thanks to which the forces on the particle in the transverse ($\hat{x},\hat{y}$) directions average out to zero 101010 With a magnetic field $B_{0}$ of the order of $10^{3}$ Gauss, the precession frequency is of the order of $10^{11}$ in the case of the SG experiment with silver atom [28].
+With the average velocity of $A_{g}$ atoms of the order of $100\,m/s$ and the size of the region of the magnetic field of the order of a few $cm$ [27], the timescale of the precession is orders of magnitude ($\sim 10^{-5}$) shorter than the time the atoms spend in the region..
+The only significant force it receives is due to the inhomogeneity in $B_{z}$, (B.2), which deflects the atom in the ±$\pm\hat{z}$ direction.
+From (B.8) and their complex conjugates, one can derive the Ehrenfest theorem for $\psi_{1}$ and $\psi_{2}$ separately,
+$\frac{d}{dt}\langle{\bf r}\rangle_{1}=\langle{\bf p}/{m}\rangle_{1}\;,\quad\frac{d}{dt}\langle{\bf p}\rangle_{1}=-\langle\nabla(\mu_{B}B_{z})\rangle_{1}\;;$ (B.10)
+$\frac{d}{dt}\langle{\bf r}\rangle_{2}=\langle{\bf p}/{m}\rangle_{2}\;,\quad\frac{d}{dt}\langle{\bf p}\rangle_{2}=+\langle\nabla(\mu_{B}B_{z})\rangle_{2}\;,$ (B.11)
+where $\langle{\bf r}\rangle_{1}\equiv\langle\psi_{1}|{\bf r}|\psi_{1}\rangle$, etc.
+Namely, for a sufficiently compact initial wave packet, $\psi({\bf r},t)$, the expectation values of ${\bf r}$ and ${\bf p}$ in the up and down components $\psi({\bf r},t)_{1,2}$ trace respectively the classical trajectories of a spin-up or spin-down particle.
+Still, even though the two subpackets might get well separated by a macroscopic distance, $\Psi$ remains in a coherent superposition of the upper and lower spin components $\psi_{1,2}$.
+Its pure-state nature can be verified, in a set-up known as the quantum eraser, by reconverging them by using a second magnetic field of opposite gradients, and studying their interferences [35].
+A variational solution of (B.8) for Gaussian split wave packets can be found in [23].

### v5

@@
 Abstract
-Quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere.
+The quantum states of a spin $\tfrac{1}{2}$ (a qubit) are parametrized by the space ${\mathbf{CP}}^{1}\sim S^{2}$, the Bloch sphere.
 A spin $j$ for a generic $j$ (a $12j+$ 1-state system) is represented instead by a point of a larger space, ${\mathbf{CP}}^{2j}$.
@@
 Vice versa, general spin states in ${\mathbf{CP}}^{2j}$ do not become classical, even at large $j$.
-We discuss these questions, by analysing the Stern-Gerlach processes, the angular-momentum composition rule, and the rotation matrix.
-Our observations help to clarify better how classical mechanics emerges from quantum mechanics in this context (e.g., with unique trajectories for a particle carrying a large spin), and to make the widespread idea that large spins somehow become classical, more precise.
+We study these questions, by analysing the Stern-Gerlach processes, the angular-momentum composition rule, and the rotation matrix.
+Our observations help to clarify better how classical mechanics emerges from quantum mechanics in this context (e.g., with unique trajectories of a particle carrying a large spin in an inhomogeneous magnetic field), and to make the widespread idea that large spins somehow become classical, more precise.
 1 Introduction
@@
 After all, the large spin limit ($j/\hbar\to\infty$) is equivalent to the $\hbar\to 0$ ($j$  fixed) limit, and, in accordance with the general idea of the semi-classical limit in quantum mechanics, a large-angular-momentum state might be expected to behave as a classical angular momentum.
-But to the best of the authors’ knowledge, it has never been clearly shown whether this indeed occurs or if it does, exactly how.
+But to the best of the authors’ knowledge, it has never been clearly shown whether this indeed occurs, or if it does, exactly how.
 It is the aim of this note to help filling in this gap.
@@
 $m\,{\dot{\bf r}}={\bf p}\;,\qquad\frac{d{\bf p}}{dt}={\bf F}=\nabla({\boldsymbol{\mu}}\cdot{\bf B})\;.$ (1.2)
-The way how such a unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [22]. The magnetic moment of a macroscopic particle is the expectation value
+The way such a unique trajectory for a classical particle emerges from quantum mechanics has been discussed in Ref. [25]. The magnetic moment of a macroscopic particle is the expectation value
 $\sum_{i}\langle\Psi|({\hat{\boldsymbol{\mu}}}_{i}+\frac{e_{i}{\hat{\boldsymbol{\ell}}}_{i}}{2m_{i}c})|\Psi\rangle={\boldsymbol{\mu}}\;,$ (1.3)
@@
 Clearly, in general, the well-known picture for a spin $\frac{1}{2}$ atom, with a doubly split wavepacket, does not generalize to a classical body (1.3) with $N\sim\mathcal{O}(10^{23})$ constituents.
-Typically, the spins are oriented along different directions, but the crucial point is that the component particles carrying them are bound in the macroscopic body in various atomic, molecular or crystalline structures.
+Typically, the spins of the component atoms and molecules are oriented along different directions, but the crucial point is that they are bound in the macroscopic body in various atomic, molecular or crystalline structures.
 The wave function of particles in bound states does not split in sub-wavepackets, in contrast to that of an isolated spin $\tfrac{1}{2}$ atom, under an inhomogeneous magnetic field.
@@
 One might wonder how the wave function of such a particle with spin $j\to\infty$ behaves, under an inhomogeneous magnetic field.
-The question is whether the three conditions recognized in [22] for the emergence of classical mechanics for a macroscopic body with a unique trajectory, reviewed in Appendix A here, are indeed sufficient.
-Or, is some extra condition, or a new unknown mechanism, needed, to suppress possible wide spreading of the wave function into many sub-packets (such as in Fig.
+The question is whether the three conditions recognized in [25] for the emergence of classical mechanics for a macroscopic body with a unique trajectory, reported here in Appendix A, are indeed sufficient.
+Or, is some extra condition, or perhaps a new unknown mechanism, needed, to suppress possible wide spreading of the wave function into many sub-packets (such as in Fig.
 1 for a small spin, $j=13/2$) under an inhomogeneous magnetic field?
-Figure 1: The distribution $|c_{k}|^{2}$ in $k$, i.e.
-in possible values of $J_{z}=m$, $-j\leq m\leq j$ ($m=-j+k$) for a spin $j=13/2$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange dots) or with $\theta=\pi/4$ (right, blue dots).
-The aim of this work is to answer these questions.
+Figure 1: The distribution in $J_{z}=-j+k$ ($k=0,1,\ldots,2j$) for a spin $j=13/2$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange dots) or with $\theta=\pi/4$ (right, blue dots).
+The wave packet of such a particle spreads into many subwavepackets à la Stern-Gerlach under an inhomogeneous magnetic field with strong gradient in the ${\hat{z}}$ direction.
+This is the central question we are going to investigate.
 We shall first discuss in Sec.
@@
 where ${\mathbf{n}}$ is a unit vector directed towards $(\theta,\phi)$ direction,
-${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;.$ (1.5)
-are to be identified with a classical angular momentum, $j\,{\mathbf{n}}$, in the large $j$ limit.
-The states (1.4) are known in the literature as coherent spin states or Bloch states [5, 10, 6, 11, 7].
+${\mathbf{n}}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\;,$ (1.5)
+or close to them, are to be identified with a classical angular momentum, $j\,{\mathbf{n}}$, in the large $j$ limit.
 Vice versa, generic states of large spin will be found to remain quantum mechanical, with large fluctuations, even in the limit, $j\to\infty$, showing that the often stated belief that a large spin (angular momentum) becomes classical, should not be taken for granted, literally.
-In subsequent sections, we are going to examine these ideas through the analysis of the Stern-Gerlach processes (Sec.
+Lest the reader is led to a misunderstanding of the content of our work, let us make the following point clear.
+The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles is quite a fruitful one both from mathematical and physical point of view.
+From the mathematical point of view, the entire theory of angular momentum can indeed be reconstructed this way [29], and it will help to recover certain formulas for large angular momentum states easily.
+It is also useful from physics point of view, as such a system may be regarded as an idealized, toy model for a more realistic macroscopic body, made of many atoms and molecules (carrying spins and orbital angular momenta).
+And this picture helps to interprete some of our findings.
+(The states of type (1.4) are known, in such a context of many-body systems, as spin coherent states, Bloch states or Glauber coherent states, depending on the author [5, 11, 6, 12, 7, 9, 15, 19, 20].
+They are all equivalent to the state of definite spin orientation (1.4), as far as the global spin property is concerned.)
+Nevertheless, our discussion is, as will be clear from an attentive reading, entirely about the quantum or classical properties of a single large spin (or angular momentum).
+We are not concerned here with the thermodynamical or other physical properties of a many-body system.
+The main question is whether a single large spin system behaves classically or remains quantum mechanical, in the limit, $j/\hbar\to\infty$.
+As will be seen in the following, the answer turns out to be quite subtle and nontrivial.
+In subsequent sections, we are going to examine these questions through the analysis of the Stern-Gerlach processes (Sec.
 3), the orbital angular momenta (Sec.
@@
 2 Space of quantum spin states and classical angular momenta
-The generic pure spin $\frac{1}{2}$ (two-state system) state is described by the wave function
+The generic pure spin $\frac{1}{2}$ (two-state system) state is described by the vector
 $|\psi\rangle=c_{1}|\!\uparrow\rangle+c_{2}|\!\downarrow\rangle\;,$ (2.1)
 where
-$(c_{1},c_{2})\sim\lambda(c_{1},c_{2})\;,\qquad\lambda\in{\mathbf{C}}/\{0\}\,$ (2.2)
+$(c_{1},c_{2})\sim\lambda(c_{1},c_{2})\;,\qquad\lambda\in{\mathbf{C}}{\backslash}{\{0\}}\,$ (2.2)
 […]
@@
 where
-$(b_{-j},b_{-j+1},\ldots,b_{j})\sim\lambda(b_{-j},b_{-j+1},\ldots,b_{j})\;,\qquad\lambda\in{\mathbf{C}}/{(0)}\;,$ (2.9)
+$(b_{-j},b_{-j+1},\ldots,b_{j})\sim\lambda(b_{-j},b_{-j+1},\ldots,b_{j})\;,\qquad\lambda\in{\mathbf{C}}{\backslash}{\{0\}}\;,$ (2.9)
 are the coordinates of the complex projective space $\mathbf{CP}^{2j}$ 333 As is well known, the manifold of pure quantum states of any $n$-state system is $\mathbf{CP}^{n-1}$ [27].
@@
 ${\hat{J}}^{2}|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\;,$ (2.15)
-have this property: its fluctuations in the components perpendicular to ${\mathbf{n}}$ become negligible in the $j\to\infty$ limit.
+(already defined in (1.4) in the Introduction) have this property: its fluctuations in the components perpendicular to ${\mathbf{n}}$ become negligible in the $j\to\infty$ limit.
 The states (2.15) are known as Bloch states or coherent spin states in the literature [5, 11, 6, 12, 7].
@@
 Now a particularly attractive picture of the quantum states (2.18) of general spin $j$, follows from the so-called stellar representation of $\mathbf{CP}^{n}$ ($n=2j$) [27]. Let the complex numbers
-$(Z_{0},Z_{1},\ldots,Z_{n})\sim\lambda\,(Z_{0},Z_{1},\ldots,Z_{n})\;,\qquad\lambda\in{\mathbf{C}}/{\{0\}}$ (2.21)
+$(Z_{0},Z_{1},\ldots,Z_{n})\sim\lambda\,(Z_{0},Z_{1},\ldots,Z_{n})\;,\qquad\lambda\in{\mathbf{C}}{\backslash}{\{0\}}$ (2.21)
 be the homogeneous coordinates of $\mathbf{CP}^{n}$, and consider the $n$-th order equation,
@@
 where $\binom{n}{k}=n!/k!(n-k)!$ are the binomial coefficients.
-The states (2.26), (2.27), coincide, apart from an overall phase factor, with the Bloch (or coherent spin) states defined in [5, 10, 6, 11, 7], and denoted as $|j,\Omega\rangle$, $|j,\{\theta,\phi\}\rangle$, etc.
+The states (2.26), (2.27), coincide, apart from an overall phase factor, with the Bloch (or spin coherent) states considered in many-body physics context [5, 11, 6, 12, 7, 9, 15, 19, 20], and denoted as $|j,\Omega\rangle$, $|j,\{\theta,\phi\}\rangle$, etc.
 In order to understand this formula from the point of view of the stellar representation of $\mathbf{CP}^{n}$, we need to translate the homogeneous coordinates of $\mathbf{CP}^{n}$ to the coefficients $\{c_{k}\}$ in (2.26) as
@@
 4, instead, are those with the minimum fluctuations, and correspond to the Bloch or spin coherent state [5, 11, 6, 12, 7].
-In the next four sections we shall illustrate how these latter states effectively behave as the classical angular momenta, $j\,{\mathbf{n}}$, in the $j\to\infty$ limit.
+In the next four sections we shall illustrate how these states effectively behave as the classical angular momenta, $j\,{\mathbf{n}}$, in the $j\to\infty$ limit.
 3 Stern-Gerlach experiments
-In this section we compare the Stern-Gerlach experiment [27] for small and large spins, with special attention to the Bloch states, (2.15), (2.26), (2.27).
+In this section we compare the Stern-Gerlach experiment [30] for small and large spins, with special attention to the states, (2.15), (2.26), (2.27).
 3.1 Spin $1/2$ atom
@@
 1 for spin $j=13/2$.
-Figure 5: The distribution $|c_{k}|^{2}$ in $k$ for possible values of $J_{z}=m$ ($m=-j+k$), for a spin $j=1\cdot 10^{3}$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange peak) or with $\theta=\pi/4$ (right, blue peak).
+Figure 5: A distribution in possible values of $J_{z}=-j+k$, for a spin $j=1\cdot 10^{3}$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange peak) or with $\theta=\pi/4$ (right, blue peak).
+The graph represents the distribution $|c_{k}|^{2}$.
 The particle starts looking like classical.
-Figure 6: The distribution $|c_{k}|^{2}$ ($J_{z}=m=-j+k$) as in Fig.
+Figure 6: The distribution in $J_{z}$ ($\equiv-j+k$) as in Fig.
 5 but for a spin $j=2\cdot 10^{5}$, for $\theta=\pi/2$ (center, orange peak) or $\theta=\pi/4$ (right, blue peak).
@@
 depending on the relative orientation of ${\mathbf{n}}_{1}$ and ${\mathbf{n}}_{2}$. Eq.(5.17) is nicely consistent with the quantum-mechanical composition law (5.1).
+5.1 Spin-orbit interactions
+As a possible variation of theme, and for completeness, let us consider a spin-orbit interaction of the form,
+$H=A\,{\mathbf{L}}\cdot{\mathbf{S}}\;$ (5.18)
+where $A$ is a constant, ${\mathbf{L}}$ and ${\mathbf{S}}$ are the orbital angular momentum and spin, respectively, of a given system.
+The problem is to understand whether and/or how such an interaction Hamiltonian reduces effectively to an interaction between two classical angular momenta, in the limit, $L/\hbar\to\infty$, $S/\hbar\to\infty$.
+By writing
+$H=\frac{A}{2}\,({\mathbf{J}}^{2}-{\mathbf{L}}^{2}-{\mathbf{S}}^{2})\;,\qquad{\mathbf{J}}={\mathbf{L}}+{\mathbf{S}}\;,$ (5.19)
+and assuming that ${\mathbf{L}}^{2}$ and ${\mathbf{S}}^{2}$ have fixed values $L(L+1)\hbar^{2}$ and $S(S+1)\hbar^{2}$ in the system under consideration, the problem reduces to that of angular momentum addition, already considered in this section.
+As such, an interaction of the form (5.18) would reduce to their classical counterpart, when the two states with angular momentum moduli $L$ and $S$ are the states of minimum fluctuations, such as $|\ell,\ell\rangle$ of (4.2) and $|j,{\mathbf{n}}\rangle$ of (2.15).
+On the other hand, for generic atomic states such conditions are not expected to be satisfied, even in the limit of large atoms.
+Thus the discussion of this section is mostly irrelevant, and the standard, quantum mechanical analysis of spin-orbit interactions are required always.
 6 Rotation matrix
@@
 3 $\sim$ Sec.
-6 we have seen how the spin coherent (Bloch) states, (2.16), behave effectively as a classical angular momentum, $j{\mathbf{n}}$, in the $j\to\infty$ limit.
+6 we have seen how the particular spin states, (2.16), behave effectively as a classical angular momentum, $j{\mathbf{n}}$, in the $j\to\infty$ limit.
 It is an entirely different story with a generic large spin $j$ state in ${\mathbf{CP}}^{2j}$,
@@
 But a macroscopic pure quantum state is possible only at exceedingly low temperature close to $T=0$: otherwise, it is a mixture.
-See [22] for more discussions.
+To show quantitatively how a macroscopically split wave packet becomes a mixed state under an environment-induced decoherence, is a complex problem, depending on the details of the environment itself, the temperature, density, pressure, kinds of environment particles, and average density and momentum distribution, the mean de Broglie wave length, etc.
+It is beyond the scope of the present work to perform such a study.
+However, a qualitative discussion about how a macroscopically split wave packet becomes a mixture under an environment-induced decoherence [34, 35, 36], but not classical, has been given in [26], in the case of a spin $\frac{1}{2}$ particle.
 To tie up possible loose ends of the discussion, note that a single, spin $\frac{1}{2}$ isolated atom in its ground state (a microscopic system) travelling in a good vacuum, is effectively a system at $T=0$.
@@
 8 Conclusion
-The observations made in this note render the concept that a spin (or angular momentum), $j$, becomes classical in the limit, $j\to\infty$, a more precise one.
-In particular, we found the identification between the states with minimum fluctuations (2.15) - the coherent spin states - with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of the SG experiments, with the angular-momentum composition rule, and with the rotation matrix.
+The observations made in this work render the concept that a spin (or angular momentum), $j$, becomes classical in the limit, $j\to\infty$, a more precise one.
+In particular, we found that the states with minimum fluctuations (2.15) and the states close to them, are to be identified with classical angular momentum vectors in such a limit, and verified its consistency through the analyses of the SG experiments, with the angular-momentum composition rule, and with the rotation matrix.
 At the same time, our analysis has revealed a subtlety in the quantum-classical correspondence in the large spin limit in general.
@@
 The reason is that in the $j\to\infty$ limit such a state necessarily involves a large space extension, in a way or other.
-A macroscopic system is in a mixed state at finite temperatures [31, 32, 33],[22].
+A macroscopic system is in a mixed state at finite temperatures [34, 35, 36, 37],[25].
 No macroscopically split pure wave packets in a SG setting survive decoherence.
 At the same time, as pointed out in [26], a macroscopic quantum state which becomes a mixture due to decoherence, remains such: a mixed quantum state.
-Decoherence in itself does not make the system considered classical, contrary to the statements often made in the literature.
-The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles turned out to be quite a fruitful one both mathematically and physically, as seen in several occasions.
-From the mathematical point of view, the entire theory of angular momentum can be reconstructed this way [26], as we recalled already, and it helps to recover certain formulas easily.
-It is also useful from physics point of view, as such a large angular momentum system may be regarded as an idealized (toy) model of a macroscopic body, made of many atoms and molecules (spins).
-Seen this way, the sharp spike of the SG projection of the state $|j,{\bf n}\rangle$ in the narrow region around $J_{z}=j\cos\theta$, Fig.
+Decoherence in itself does not render the system involved classical, contrary to such an idea sometimes expressed in the literature [35, 37].
+The idea that a large spin is made of many spin $\tfrac{1}{2}$ particles turned out to be quite a useful one both mathematically and physically, as we saw in several occasions.
+The entire theory of angular momentum can indeed be reconstructed this way as shown by Schwinger [29], and it helps to understand certain formulas easily.
+It is also relevant from the physics point of view, as such a large angular momentum system may be regarded as an idealized toy-model version of a macroscopic body, made of many atoms and molecules (spins).
+Seen this way, the sharp spike of the SG projection of the state $|j,{\bf n}\rangle$ in the narrow region around $J_{z}=j\cos\theta$ we have observed, such as in Fig.
 5, Fig.
-6, can be understood as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $J_{z}$.
+6, can be interpreted as a results of a large number of microscopic states of atoms (spins) accumulating to give such a value of $J_{z}$.
 [^8] The suppression of the fluctuations of angular momentum components we observed in this note in the $j\to\infty$ limit, is somewhat analogous to the “$\sqrt{n}$ law” invoked by Schrödinger [38], to explain the general exactness (classicality) of laws of macroscopic world based on statistical mechanics, although, here, the relevant fluctuations are quantum-mechanical ones.
 This being so, let us not forget that the deepest insight of Schrödinger in his book was that many biological processes such as mutation, reproduction, and inheritance, involve quantum mechanics in essential ways.
-Nevertheless, logically our discussion is about the quantum and classical properties of a single large spin, and is not about the thermodynamical or statistical properties of many body systems, as in [6, 7], for instance.
-And in spite of inevitable partial overlaps with numerous observations made in the literature [1]-[21], the ideas around the proposal (2.16) exposed in Sec.
-2, Sec.
-3, Sec.
+Nevertheless, our discussion is really about the quantum and classical properties of a single large spin, and do not concern the thermodynamical, statistical or other physical properties of realistic many-body systems, as in [6, 7].
+And in spite of inevitable partial overlaps with numerous observations made in the literature [1]-[24], our careful observations about how the particular class of large momentum (spin) states get classical in the limit, $j\to\infty$, independently of the temperature, whereas a generic large $j$ states remain quantum mechanical in the same limit, as illustrated in Sec.
+2 - Sec.
 5, and Sec.